### v2

# A Deep Learning-Based Supervised Transfer Learning Framework for DOA Estimation with Array Imperfections

Abstract
In practical scenarios, processes such as sensor design, manufacturing, and installation will introduce certain errors. Furthermore, mutual interference occurs when the sensors receive signals. These defects in array systems are referred to as array imperfections, which can significantly degrade the performance of Direction of Arrival (DOA) estimation. In this study, we propose a deep-learning based transfer learning approach, which effectively mitigates the degradation of deep-learning based DOA estimation performance caused by array imperfections.
In the proposed approach, we highlight three major contributions. First, we propose a Vision Transformer (ViT) based method for DOA estimation, which achieves excellent performance in scenarios with low signal-to-noise ratios (SNR) and limited snapshots. Second, we introduce a transfer learning framework that extends deep learning models from ideal simulation scenarios to complex real-world scenarios with array imperfections. By leveraging prior knowledge from ideal simulation data, the proposed transfer learning framework significantly improves deep learning-based DOA estimation performance in the presence of array imperfections, without the need for extensive real-world data. Finally, we incorporate visualization and evaluation metrics to assess the performance of DOA estimation algorithms, which allow for a more thorough evaluation of algorithms and further validate the proposed method. Our code can be accessed at https://github.com/zzb-nice/DOA_est_Master.
Index Terms:
Direction of Arrival (DOA) estimation, Domain Adaptive, Vision Transformer, Data-driven Methods, Deep learning
I INTRODUCTION
Direction of Arrival (DOA) estimation finds extensive applications across various engineering fields, including wireless communications, astronomical observation, radar, and sonar [2, 3, 4, 5, 6], where accurately determining the direction of incoming signals is crucial for efficient operation. Today, as radar technology becomes increasingly accessible for civilian use, DOA estimation algorithms have also been applied in emerging areas such as autonomous driving and drone operating system [7, 8, 9, 10].
With the continuous development of engineering technologies, practical scenarios are placing higher demands on DOA estimation algorithms. For example, in autonomous driving scenarios, there is an increasing need for DOA estimation of fast-moving small targets [7, 11]. In non-cooperative target localization scenarios, it is essential to achieve rapid and precise DOA estimation [12, 13]. These practical scenarios generally require accurate DOA estimation results under conditions of low SNR and small snapshots, along with high real-time performance. Therefore, there is an urgent need for a new algorithm with relatively low computational complexity, one that can achieve strong performance in ideal scenarios while also demonstrate improved performance in low SNR and small snapshots scenarios.
Over the past few decades, extensive research has been conducted on DOA estimation, leading to the development of numerous algorithms [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 1, 29, 30, 31, 32, 33, 34, 35]. These methods can be broadly categorized into model-driven and data-driven approaches. Among them, model-driven methods have been more extensively studied due to their well-established theoretical foundations. The core principle of model-driven DOA estimation is to establish a mapping between the directions of incoming signals and the received array data, and then use a series of techniques to accurately invert this mapping based on the received data. Based on different estimation strategies, model-driven approaches can be further classified into several major categories, including beamforming methods [14], subspace methods [15, 16, 17, 18, 19], maximum likelihood methods [20, 21, 22], and compressed sensing (sparsity-inducing) methods [36, 23, 24, 25].
A leading and well-respected family of model-driven approaches is the subspace method, which achieves satisfactory performance in scenarios with high signal-to-noise ratio (SNR) and large snapshots. It maintains relatively low computational complexity when a limited number of array elements is used. However, its performance deteriorates significantly in practical scenarios with low SNR and small snapshots. Moreover, subspace-based methods struggle with color noise, and require the estimated number of signal sources to be smaller than the number of array elements, which limits their applicability. Compressive sensing-based and maximum likelihood-based methods have achieved certain improvements but are limited by real-time performance constraints. In summary, model-driven methods for DOA estimation have reached a performance bottleneck in practical scenarios.
In recent years, the rapid advancement of computational hardware [37] has significantly accelerated the development of deep learning technologies. As data-driven technologies, machine learning and deep learning methods have gained increasing attention in the field of DOA estimation. These methods build complex mapping relationships from input to output based on large-scale datasets. Unlike model-driven methods that rely on predefined signal models, data-driven methods generally do not impose prior assumptions on the input and output. This makes them particularly suitable for practical scenarios where the signal model may be incomplete or inaccurate. Furthermore, during practical implementation, data-driven methods demonstrate better real-time performance due to their fast inference speed. As a result, data-driven methods are considered promising tools for achieving DOA estimation in practical scenarios.
Deep learning methods were first introduced to address the problem of DOA estimation in the 1990s [26, 27]. Following a considerable period of minimal utilization, these approaches have witnessed a substantial resurgence in DOA estimation research during recent years. The investigations conducted in [28, 1] pioneered the implementation of Deep Neural Network (DNN) and Convolutional Neural Network (CNN) for DOA estimation. Subsequently, the method proposed in [29, 30] advanced the field by first partitioning the incoming signal into multiple sub-regions through a network architecture, followed by the deployment of parallel networks across the sub-regions.
Additionally, some methods aim to combine traditional DOA estimation techniques with deep learning methods. Due to their inherent flexibility, deep learning methods can effectively establish diverse mapping relationships. As a result, deep learning techniques are not necessarily limited to direct mapping to the DOA values but can also be applied to map various variables, such as noise-free covariance matrices, signal subspaces, angular intervals of incident signals, and so on. This flexibility makes them well-suited for integration with traditional methods such as subspace techniques and compressed sensing. In 2022, the method proposed in [31, 32] integrated subspace-based methods with deep neural networks. Specifically, a CNN was employed to reconstruct a noise-free covariance matrix, after which a classical subspace-based approach was applied for DOA estimation. Similarly, the approaches in [33, 34] combined deep learning with compressed sensing algorithms. In these works, a DNN is utilized to the extend the covariance matrix to a higher-dimensional space or to compute prior knowledge for the compressed sensing algorithm. Then a sparse reconstruction methods is applied to estimate the incident angles. Furthermore, in 2024, Shmuel et al. [35] introduced SubspaceNet, which directly treated the Root MUSIC algorithm as part of the network. Unlike previous work, SubspaceNet utilized the differentiability of polynomial equation root-finding and matrix decomposition, which allowed for the backpropagation of loss from the Root MUSIC algorithm.
Nevertheless, most of the model-driven and data-driven methods mentioned above do not account for array imperfections in practical systems. The existence of array position biases, amplitude and phase inconsistencies, mutual coupling effects [38], and other array imperfections can affect the DOA estimation accuracy of models. In severe cases, these imperfections can significantly degrade the model’s DOA estimation performance or even cause its DOA estimation to fail.
A series of active calibration [39, 40, 41, 42, 43] and self-calibration [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54] methods have been applied to compensate for array imperfections in practical scenarios. Among them, active calibration methods require the setup of known signal sources with predefined DOA directions, which are specifically employed for the compensation of array imperfections. In contrast, self-calibration methods do not require additional auxiliary signal sources and can dynamically calibrate the array imperfections during array system applications, offering significant practical advantages. However, active calibration methods benefit from utilizing known DOA values for array imperfection estimation, leading to higher accuracy. In contrast, self-calibration methods generally estimate array imperfections simultaneously with DOA values, often relying on initialization and struggling to handle more severe array imperfections. As noted in [44], most existing self-calibration methods can perform well only under relatively ideal signal conditions and are ineffective in practical scenarios with low SNR, small snapshots, and spatially proximate sources——precisely the challenges that many DOA estimation algorithms face.
Moreover, existing calibration methods are constrained by numerous limitations. Due to the significant difficulty in directly modeling and analyzing all array imperfections, most existing calibration algorithms can focus on only one of the array imperfections, such as array position biases [39, 42, 43], amplitude-phase inconsistencies [45, 46, 47, 48, 49] or mutual coupling [50, 51, 52, 53, 54]. Only a few studies have explored how to address scenarios involving multiple array imperfections [44, 40, 41]. Some methods are even restricted to calibrating specific array structures such as uniform linearity/circularity array [54, 48, 49, 53], or impose additional assumptions to simplify the model [44]. Therefore, how to effectively address scenarios with multiple array imperfections remains an issue that requires further exploration.
In [29], Liu et al. mentioned that deep learning algorithms exhibit good adaptability to array imperfections, but this adaptability fails in practical scenarios with low SNR and small snapshots. Based on the simulation results, it can be observed that in scenarios with low SNR and small snapshots, deep learning models remain sensitive to array imperfections, requiring the application of calibration models to compensate for array imperfections. However, due to the inherent flexibility of deep learning, it may be possible to directly construct a transfer learning algorithm [55] that transitions from an ideal scenario without array imperfections to a practical scenario with array imperfections, thereby enhancing the model’s adaptability to practical scenarios.
Nonetheless, there is currently a lack of research on the application of transfer learning methods in the field of DOA estimation. To the best of our knowledge, only a few papers [56, 57, 58] have introduced transfer learning methods into DOA estimation. In [56], Wu et al. migrated a network designed for DOA estimation with multiple targets to one for fewer targets, with the primary goal of reducing the computational cost of training. Similarly, in [58], Labbaf et al. migrated a network designed for single-target DOA estimation to multiple targets. In [57], Zhang et al. applied transfer learning to enhance the network’s adaptability to array imperfections, but they only fine-tuned the network from ideal data to array imperfections-affected data, without incorporating more advanced transfer learning techniques.
In this paper, we propose a deep learning-based supervised transfer learning framework for DOA estimation, which aims to mitigate the performance degradation of data-driven approaches in practical scenarios with various array imperfections. In our approach, the data is divided into source and target domains. The source domain data consists of ideal data without array imperfections, while the target domain data represents practical scenarios containing array imperfections. Firstly, we train a DOA estimation model using the source domain data. A transfer learning method is then employed to align the features between the source and target domains, thereby enhancing the deep learning model’s ability to adapt to array imperfections and improving performance in scenarios with low SNR, small snapshots, and multiple array imperfections. Moreover, in a variety of scenarios, extensive experiments have been conducted to validate the effectiveness of the proposed model. The contributions of our work can be outlined as follows:
  1. 1.
    We propose a novel DOA estimation model based on Vision Transformer, which demonstrates exceptional performance under challenging conditions, such as low SNR and small snapshot scenarios.
  2. 2.
    Due to the presence of array imperfections, the data distributions of the source and target domains differ significantly, leading to substantial performance degradation when the model is deployed in practical scenarios. To address this issue, we introduce a transfer learning algorithm to align the features between the source and target domains, enhancing the model’s performance in practical scenarios.
  3. 3.
    Via extensive simulations, we compare the proposed method with existing approaches across multiple evaluation metrics and demonstrate the superiority of our method in terms of DOA estimation accuracy and robustness.
II Problem Formulation
Figure 1: The framework of DOA-ViT and transfer learning approach proposed for array imperfections.
In this work, two models of array receiving signals are considered. The first is the ideal array signal model, where the signal received by each sensor is represented as $Be^{j\varphi}s(t)$. where $B$ denotes the sensor’s gain for the incoming signal, and $e^{j\varphi}$ accounts for the phase shift caused by the time delay of the signal propagation between the sensors. This idealized model aligns with the standard assumptions used in most DOA estimation methods.
The second signal model incorporates array imperfections to reflect real-world conditions. It considers factors such as gain and phase inconsistencies, sensor position biases, and mutual coupling. These imperfections introduce additional “array imperfection noise”, which affects the received signal and captures the complexities of practical scenarios.
II-A Ideal Signal Model
Firstly, the ideal array reception model is constructed. Consider a Uniform Linear Array (ULA) or Uniform Circular Array (UCA) consisting of $M$ antennas, receiving $K$ narrowband non-coherent far-field signals impinging from different DOAs. The array received signal model can be expressed in the following manner:
$\boldsymbol{y}(t)$ $=\sum_{k=1}^{K}\boldsymbol{a}\left(\boldsymbol{p},\theta_{k}\right)s_{k}(t)+\boldsymbol{n}(t)$ (1)
$=\boldsymbol{A}(\boldsymbol{p},\boldsymbol{\theta})\boldsymbol{s}(t)+\boldsymbol{n}(t),t=1,\ldots,T.$
where $\boldsymbol{\theta}=\left[\theta_{1},\ldots,\theta_{K}\right]$ represents the directions of arrival, $\boldsymbol{A}(\boldsymbol{p},\boldsymbol{\theta})=\left[\boldsymbol{a}\left(\boldsymbol{p},\theta_{1}\right),\ldots,\boldsymbol{a}\left(\boldsymbol{p},\theta_{K}\right)\right]$ denotes the $M\times K$ array manifold matrix, $t=1,\ldots,T$ is the time index, and $T$ is the number of received snapshots. $\boldsymbol{s}(t)={{[{s_{1}}(t),\ldots,{s_{K}}(t)]}^{T}}$ and $\boldsymbol{n}(t)$ denote the transmitted signal and the noise vector at time index $t$, respectively. $\boldsymbol{s}(t)$ is Gaussian distributed in both its real and imaginary parts, and $\boldsymbol{n}(t)$ is additive Gaussian noise, with no correlation between any of the signals or noise. For a ULA, $\boldsymbol{a}(\boldsymbol{p},\theta_{k})$ can be expressed as:
$\boldsymbol{a}\left(\boldsymbol{p},\theta_{k}\right)=\left[1,e^{-j\frac{2\pi d}{\lambda}\sin\left(\theta_{k}\right)},\ldots,e^{-j\frac{2\pi d}{\lambda}(M-1)\sin\left(\theta_{k}\right)}\right]^{T}$ (2)
where $\lambda=c/f$ denotes the wavelength of the incident signals, $d$ is the distance between the array elements, while $c$ and $f$ represent the propagation speed and the frequency of the incident signals, respectively. Similarly, for a UCA, $\boldsymbol{a}(\boldsymbol{p},\phi_{k},\theta_{k})$ can be represented in the following form:
$\boldsymbol{a}\left(\boldsymbol{p},\phi_{k},\theta_{k}\right)=$ $\left[e^{-j\left(2\pi/\lambda\right)\cos(\phi_{1}-\theta_{k})\sin(\phi_{k})},\dots,\right.$ (3)
$\left.e^{-j\left(2\pi/\lambda\right)\cos(\phi_{M}-\theta_{k})\sin(\phi_{k})}\right]^{T}$
where $\phi_{m}=\frac{2\pi m}{M}$ denotes the azimuth of the $m$th sensor of the UCA. Based on the received signal vector discussed above, the spatial covariance matrix (SCM) is computed as follows:
[…]
where $\mathbb{E}\{\cdot\}$ denotes the expectation operator, ${{\boldsymbol{R}}_{s}}=\mathbb{E}\left\{\boldsymbol{s}(t){{\boldsymbol{s}}^{\mathrm{H}}}(t)\right\}$ represents the covariance matrix of incident signals and $\sigma_{N}^{2}$ stands for the noise power. Depending on the statistical degrees of freedom of SCM, $K\leq M$ different DOAs can be estimated precisely. However, since the number of snapshots is limited in practice, the SCM is replaced by the sampled SCM calculated via:
$\widetilde{\boldsymbol{R}}_{y}=\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{y}(t)\boldsymbol{y}^{\mathrm{H}}(t)$ (5)
$\widetilde{\boldsymbol{R}}_{y}$ is an unbiased estimate of ${\boldsymbol{R}}_{y}$ calculated by the received signal vector. Most existing DOA estimation algorithms focus on deriving the source DOAs $\boldsymbol{\theta}=[{{\theta}_{1}},\ldots,{{\theta}_{K}}]$ from the sampled SCM $\widetilde{\boldsymbol{R}}_{y}$, which has been proven to be the most effective approach.
II-B Signal Model with Array Imperfection
In practical applications, the array signal reception model is often not as ideal as its theoretical counterpart. Position biases may be introduced during the arrangement of array elements. Amplitude and phase inconsistencies may be caused by defects in the design and manufacturing of sensors. And mutual coupling effects may be induced during the signal reception process. In [29], Liu et al. established a simplified model to show how these three array imperfections affect the signal model in ULA, and introduced a hyperparameter $\rho\in[0,1]$ to succinctly control the intensity of these array imperfections.
$\boldsymbol{e}_{\text{pos}}&=\rho\times\left[0,-0.2,\dots,-0.2,0.2,0.2,\dots,0.2\right]^{T}\times d\\ \boldsymbol{e}_{\text{gain}}&=\rho\times\left[0,0.2,\dots,0.2,-0.2,\dots,-0.2\right]^{T}\\ \boldsymbol{e}_{\text{phase}}&=\rho\times\left[0,-30^{\circ},\dots,-30^{\circ},30^{\circ},\dots,30^{\circ}\right]^{T}\\ \boldsymbol{e}_{\text{mc}}&=\rho\times\left[\gamma^{0},\gamma^{1},\dots,\gamma^{M-1}\right]^{T}\end{aligned}$
where $\gamma=0.3e^{j60^{\circ}}$ is the mutual coupling coefficient between adjacent sensors. The term $\boldsymbol{e}_{\text{pos}}$ represents the positional bias of each array element, while $\boldsymbol{e}_{\text{gain}}$ and $\boldsymbol{e}_{\text{phase}}$ denote the amplitude and phase biases, respectively. $\boldsymbol{e}_{\text{mc}}$ models the magnitude of the mutual coupling effect for a ULA, where the strength of the mutual coupling is proportional to the square of the distance. The matrix $\boldsymbol{E}_{\text{mc}}$, denoted as a Toeplitz matrix with parameter vector $\boldsymbol{e}_{\text{mc}}$ [59], can be employed to model the mutual coupling across the entire array:
$\small\boldsymbol{E}_{\text{mc}}=\begin{bmatrix}\gamma^{0}&\gamma^{1}&\cdots&\gamma^{i}&\gamma^{i+1}&\cdots&\gamma^{M-1}\\ \gamma^{1}&\gamma^{0}&\gamma^{1}&\cdots&\gamma^{i}&\cdots&\vdots\\ \vdots&\gamma^{1}&\gamma^{0}&\ddots&\cdots&\ddots&\vdots\\ \gamma^{i}&\cdots&\ddots&\ddots&\ddots&\ddots&\gamma^{i}\\ \gamma^{i+1}&\ddots&\cdots&\ddots&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&\cdots&\ddots&\gamma^{0}&\gamma^{1}\\ \gamma^{M-1}&\cdots&\gamma^{i+1}&\gamma^{i}&\cdots&\gamma^{1}&\gamma^{0}\end{bmatrix}$ (7)
Similarly, for a UCA, the $\boldsymbol{e}_{\text{mc}}$ is exponentially related to the inter-element spacing. The position bias $\boldsymbol{e}_{\text{pos}}$ can be extended to both the $x$- and $y$-axes, while $\boldsymbol{e}_{\text{gain}}$ and $\boldsymbol{e}_{\text{phase}}$ remain unchanged. The steering vector affected by all three array imperfections can be expressed as:
$\boldsymbol{a}(\theta,\boldsymbol{e})&=\left(\boldsymbol{I}_{M}+\delta_{mc}\boldsymbol{E}_{mc}\right)\times\left(\boldsymbol{I}_{M}+\operatorname{Diag}\left(\delta_{gain}\boldsymbol{e}_{gain}\right)\right)\\ &\times\operatorname{Diag}\left(\exp\left(j\delta_{\text{phase}}\boldsymbol{e}_{\text{phase}}\right)\right)\times\boldsymbol{a}\left(\theta,\boldsymbol{p}+\delta_{pos}\boldsymbol{e}_{pos}\right)\end{aligned}$
where $\delta(\cdot)$ is used to denote specific array imperfections, $I_{M}$ represents the $M\times M$ identity matrix, while $\operatorname{Diag}(\cdot)$ denotes a diagonal matrix formed from a given vector. And $\boldsymbol{a}\left(\theta,\boldsymbol{p}+\boldsymbol{e}_{pos}\right)$ refers to the array steering vector obtained based on the position biases:
$\boldsymbol{a}\left(\theta_{k},\boldsymbol{p}+\boldsymbol{e}_{pos}\right)=$ $[e^{-j\frac{2\pi}{\lambda}(0d+e_{\text{pos}}^{0})\sin(\theta_{k})},e^{-j\frac{2\pi}{\lambda}(1d+e_{\text{pos}}^{1})\sin(\theta_{k})},$ (9)
$\ldots,e^{-j\frac{2\pi}{\lambda}((N-1)d+e_{\text{pos}}^{N-1})\sin(\theta_{k})}]^{T}$
where $e_{\text{pos}}^{i}$ represents the position biases of the $i$th antenna. Therefore, the final result $\boldsymbol{a}(\theta,\boldsymbol{e})$ in Equation (8) represents the steering vector influenced by the three types of array imperfections. Therefore, the imperfection steering matrix $\boldsymbol{A}$ can be expressed as:
$\boldsymbol{A}(\boldsymbol{\theta},\boldsymbol{e})=\left[\boldsymbol{a}(\theta_{1},\boldsymbol{e}),\boldsymbol{a}(\theta_{2},\boldsymbol{e}),\ldots,\boldsymbol{a}(\theta_{K},\boldsymbol{e})\right]$ (10)
By substituting $\boldsymbol{A}(\boldsymbol{\theta},\boldsymbol{e})$ for the steering matrix $\boldsymbol{A}(\boldsymbol{p},\boldsymbol{\theta})$ in Section II-A, a more accurate simulation of the array signal reception model in practical scenarios can be achieved.
III NETWORKS FOR DOA ESTIMATION
Figure 2: The overrall structure of DOA-ViT and the processing procedure of input covariance matrix.
Given the remarkable performance of the Vision Transformer (ViT) model across various domains, its strong nonlinear representation capability and efficient feature learning ability are expected to enhance subsequent transfer learning. Therefore, we adopt ViT as our baseline model. The overall architecture of the ViT model designed for DOA estimation, referred to as DOA-ViT, is illustrated in Figure 2. In terms of model design, we adhere to the original Vision Transformer framework [60] as closely as possible to ensure consistency and leverage its proven architecture.
III-A Covariance Matrix Embedding
Since the transformer encoder can only operate on a set of input vectors, it is necessary to encode the input covariance matrix into a set of vectors.
To achieve this, the split and vectorization operation is employed, which is detailed in (11) and (12). The sampled SCM $\widetilde{\boldsymbol{R}}_{y}$ is partitioned column-wise into $M$ smaller blocks and subsequently each block is flattened into a vector of size $2M$, resulting in $M$ vectors. Finally, a linear layer is utilized to encode each input vector, generating a matrix of dimensions $M\times D$, which represents $M$ vectors, each of dimension $D$.
$\widetilde{\boldsymbol{R}}_{y}$ $=\left[\boldsymbol{r}_{1},\boldsymbol{r}_{2},\boldsymbol{r}_{3},\ldots,\boldsymbol{r}_{M}\right],\;\small{\widetilde{\boldsymbol{R}}_{y}\in\mathbb{C}^{M\times M},r_{i}\in\mathbb{C}^{M}}$ (11)
$\widetilde{\boldsymbol{r}}_{i}$ $=\left[\text{real}(\boldsymbol{r}_{i});\text{imag}(\boldsymbol{r}_{i})\right],\quad\widetilde{\boldsymbol{r}}\in\mathbb{R}^{2M}$ (12)
$\boldsymbol{x}_{i}$ $=\boldsymbol{E}\widetilde{\boldsymbol{r}}_{i},\quad\boldsymbol{E}\in\mathbb{R}^{D\times 2M}$ (13)
$\boldsymbol{X}_{\text{input}}$ $=\left[doa\;token,\boldsymbol{x}_{1},\boldsymbol{x}_{2},\ldots,\boldsymbol{x}_{M}\right]$ (14)
The formulation of Covariance Matrix Embedding procedure is presented in Eqs. (11) to (14). Notably, similar to the widely used transformer models, a learnable $D$-dimensional vector, called the DOA estimation token, is added to the first column of the embeded matrix. Since the transformer encoder captures global dependencies through self-attention, the DOA estimation token effectively integrates information from the entire sequence and is subsequently used in the final task, such as classification and regression. This approach simplifies model design and enables the transformer model to directly leverage global context for task prediction.
III-B Position Embedding
Transformer models inherently lack information about the order of input vectors. In other words, even if the positions of input vectors are randomly shuffled, the transformer encoder produces identical outputs for each input vector, as it treats the input vectors as an unordered set. To address this limitation, positional embeddings are introduced to encode the relative or absolute positions of the input vectors. Specifically, a set of learnable vectors, known as positional embeddings, is added directly to the input vectors as additional positional encoding information:
$\boldsymbol{z}_{\text{input }}$ $=\left[\boldsymbol{x}_{0}+\boldsymbol{e}_{0},\boldsymbol{x}_{1}+\boldsymbol{e}_{1},\cdots,\boldsymbol{x}_{M}+\boldsymbol{e}_{M}\right]$ (15)
$=\boldsymbol{X}_{\text{input }}+\boldsymbol{E}_{\text{pos }},\quad\boldsymbol{E}_{\text{pos }}\in\mathbb{R}^{D\times(M+1)}$
By adding position embeddings to the input representations, the model can incorporate positional information into its computations, ensuring that the order of the input vectors influences the final representation. This mechanism is essential for tasks where understanding the sequential arrangement of input vectors significantly impacts the model’s performance, such as DOA estimation.
III-C Transformer Encoder
Figure 3: The framework of transfer learning approach proposed for array imperfections.
The Multi-Head Self-Attention (MHSA) mechanism is a key component of the Transformer Encoder, which is designed to enhance the model’s capacity to capture complex relationships within sequences.
Given an input $\boldsymbol{Z}\in\mathbb{R}^{D\times(M+1)}$, where $D$ is the feature dimension and $M+1$ is the sequence length, the MHSA operates are summarized as follow.
Linear Projections: The input is linearly projected into queries $\boldsymbol{Q}$, keys $\boldsymbol{K}$, and values $\boldsymbol{V}$ for each attention head $i$:
$\boldsymbol{Q}_{i}=\boldsymbol{W}_{i}^{Q}\boldsymbol{Z},\quad\boldsymbol{K}_{i}=\boldsymbol{W}_{i}^{K}\boldsymbol{Z},\quad\boldsymbol{V}_{i}=\boldsymbol{W}_{i}^{V}\boldsymbol{Z},\quad i=1,\dots,h$ (16)
where $\boldsymbol{W}_{i}^{Q},\boldsymbol{W}_{i}^{K},\boldsymbol{W}_{i}^{V}\in\mathbb{R}^{d_{k}\times D}$ are model parameters, and $d_{k}=D/h$ is the dimension of each head.
Scaled Dot-Product Attention: Each head computes attention as:
$\text{head}_{i}=\text{softmax}\left(\frac{\boldsymbol{Q}_{i}\boldsymbol{K}_{i}^{T}}{\sqrt{d_{k}}}\right)\boldsymbol{V}_{i}$ (17)
Concatenation and Output: The outputs of all $h$ heads are concatenated and projected to produce the final output:
$\text{MHSA}(\boldsymbol{Z})=\boldsymbol{W}^{O}\text{Concat}(\text{head}_{1},\dots,\text{head}_{h})$ (18)
where $W^{O}\in\mathbb{R}^{D\times hd_{k}}$ is the output projection matrix. As illustrated in Fig. 2, the Transformer Encoder, consisting of layer normalization, MHSA and an MLP, enables the DOA-ViT model to effectively capture and integrate global information.
III-D DOA Estimation Head
The DOA Estimation Head serves as the final layer for DOA estimation. In this module, the vector corresponding to the DOA estimation token is extracted; then a linear layer is directly applied to regress the target DOA values. The loss function between the predicted DOA values and the true DOA is expressed as:
$l(\boldsymbol{\widetilde{\boldsymbol{R}}},\boldsymbol{\theta};\boldsymbol{\psi})=\min_{\boldsymbol{P}\in\mathcal{P}}\left(\frac{1}{K}\left\|\boldsymbol{\theta}-\boldsymbol{P}\hat{\boldsymbol{\theta}}(\widetilde{\boldsymbol{R}};\boldsymbol{\psi})\right\|^{2}\right)^{\frac{1}{2}}$ (19)
where $\boldsymbol{\psi}$ represents all the parameters of the model and $\hat{\boldsymbol{\theta}}$ is the estimation result of the DOA-ViT model, which includes $K$ DOAs. $\boldsymbol{\theta}$ corresponds to the true DOAs, and the set $\mathcal{P}$ encompasses all $M\times M$ permutation matrices, defined as binary matrices where each row and column contains exactly one non-zero element. The loss function for a dataset $\mathcal{D}$ is described in the following manner:
$\mathcal{L}_{\mathcal{D}}(\boldsymbol{\psi})=\frac{1}{|\mathcal{D}|}\sum_{(\boldsymbol{\theta}^{(j)},\widetilde{\boldsymbol{R}}^{(j)})\in\mathcal{D}}l(\widetilde{\boldsymbol{R}}^{(j)},\boldsymbol{\theta}^{(j)};\boldsymbol{\psi})$ (20)
where $|\mathcal{D}|$ denotes the number of samples in the simulated dataset and $\widetilde{\boldsymbol{R}}^{(j)}$ is the $j$th sampled SCM in the dataset. It is worth noting that our proposed method can be easily extended to arbitrary array configurations and 2D (Two-Dimensional) DOA estimation. If one aims to implement 2D DOA estimation, the DOA Estimation Head can be directly extended to predict the values of $\theta$ and $\phi$. The corresponding loss function for this approach is derived in the following form:
$\begin{split}l(\boldsymbol{\widetilde{\boldsymbol{R}}},\boldsymbol{\theta},\boldsymbol{\phi};\boldsymbol{\psi})=&\min_{\boldsymbol{P}\in\mathcal{P}}\left(\frac{1}{2K}\left\|\boldsymbol{\theta}-\boldsymbol{P}\hat{\boldsymbol{\theta}}(\widetilde{\boldsymbol{R}};\boldsymbol{\psi})\right\|^{2}\right.\\ &\quad\left.+\frac{1}{2K}\left\|\boldsymbol{\phi}-\boldsymbol{P}\hat{\boldsymbol{\phi}}(\widetilde{\boldsymbol{R}};\boldsymbol{\psi})\right\|^{2}\right)^{\frac{1}{2}}\end{split}$ (21)
IV Transfer Learning Approach
By leveraging large-scale datasets for training, deep neural networks have achieved superior performance in DOA estimation. However, due to the statistical differences in the sampled SCM between simulated data and real-world data with array imperfections, the model’s performance in practical scenarios is adversely affected.
Inspired by recent advancements in transfer learning [61, 62], and in order to address the significant performance degradation of deep learning models due to array imperfections in practical scenarios, we propose a supervised transfer learning approach, which is designed to enhance the adaptability of deep learning methods to array imperfections.
IV-A Goal of Transfer Learning
Transfer learning typically involves two tasks: the source task and the target task. The source task usually has abundant data available to effectively train the model, whereas the target task has a limited amount of data. Building on machine learning methods, transfer learning allows prior knowledge from the source task to be applied to the target task, even though there are different data distributions between the source and target tasks [55].
IV-B Source and Target Domains Modeling
In our transfer learning approach, two distinct scenarios are considered. The first is the ideal array signal reception scenario, which serves as a scenario that most deep learning models focus on. In this scenario, training data can typically be generated in large quantities through simulation programs, resulting in ample data availability at a relatively low cost. Consequently, deep learning algorithms can be effectively trained and validated based on the extensive dataset, ultimately achieving satisfactory performance.
The second scenario pertains to practical applications, where array imperfections in the antenna system are present. In this case, data should be obtained through actual measurements, leading to higher costs and a relatively limited dataset.
The ideal scenario data is treated as source domain data, and the practical scenario data as target domain data. Therefore, how to effectively leverage the prior knowledge acquired from a model trained on a substantial amount of simulation data to enhance model’s performance in a data-scarce or even data-absent practical setting becomes a valuable area for exploration.
IV-C Transfer learning Framework
Assume that we have already trained a model with weight parameters $\boldsymbol{\psi}_{s}$ that performs well on simulated data. This model is capable of effectively extracting features and performing DOA estimation in the source domain. However, due to the presence of array imperfections, substantial statistical discrepancies arise between the data distributions of the source and target domains. Consequently, when applied to the target domain, the model with parameters $\boldsymbol{\psi}_{s}$ extracts features that are significantly influenced by these statistical discrepancies, thereby deteriorating DOA estimation performance.
To mitigate this issue, the objective of this work is to train a feature extractor for the target domain that is designed to learn domain-invariant features, which remain unaffected by domain shifts induced by array imperfections, such as element position biases, amplitude-phase inconsistencies and mutual coupling effects that distort the statistical properties of the sampled SCM. Only the essential information relevant to the DOA estimation task is expected to be retained in domain-invariant features, thereby enhancing the model’s generalization ability in practical scenarios. The framework of our transfer learning framework is illustrated in Fig. II, and we will provide a detailed introduction in the following.
In our transfer learning approach, we remove the DOA estimation Head from the DOA-ViT model and treat the remaining structure as a feature extractor, denoted as $\mathcal{F}$, which allows us to extract features from both the source and target domains:
$\boldsymbol{z}_{s,i}$ $=\mathcal{F}_{s}(\widetilde{\boldsymbol{R}}_{s}^{(i)};\boldsymbol{\psi}_{s})$ (22)
$\boldsymbol{z}_{t,i}$ $=\mathcal{F}_{t}(\widetilde{\boldsymbol{R}}_{t}^{(i)};\boldsymbol{\psi}_{t})$
where $\mathcal{F}_{s}$ and $\mathcal{F}_{t}$ represent the feature extractors for the source and target domains, respectively, while $\boldsymbol{\psi}_{s}$ and $\boldsymbol{\psi}_{t}$ represent the model parameters for the source and target domain models. $\boldsymbol{z}_{s,i}$ represents the features extracted from the $i$-th source domain input, and $\boldsymbol{z}_{t,i}$ denotes the features extracted from the $i$-th target domain input.
In practical scenarios, a dataset can be constructed by collecting a series of sampled SCMs along with their corresponding DOA values. Base on the true DOA values, a simulation data generator is employed to generate corresponding ideal covariance matrices that are unaffected by array imperfections, which serve as the source domain data to be aligned. As illustrated in Fig. II, during the transfer learning approach, data in the source domain is fed into $\mathcal{F}_{s}$ with frozen pre-trained weights to ensure the effectiveness of feature extraction from the source domain, while data in the target domain is fed into a trainable feature extraction module $\mathcal{F}_{t}$. In this way, by aligning the features $\boldsymbol{z}_{t,i}$ to $\boldsymbol{z}_{s,i}$, the target domain feature extractor $\mathcal{F}_{t}$ can be guided to extract features that can be directly used for DOA estimation, thereby mitigating the impact of array imperfections. Based on the feature alignment procedure described above, two loss functions are proposed for feature alignment as follows.
MSE Loss: A mean squared error is directly applied between the features of the source and target domains to achieve the alignment of target domain features with source domain features.
$\mathcal{L}_{\text{mse}}(\boldsymbol{z}_{i}^{S\leftrightarrow T})=\|\boldsymbol{z}_{s,i}-\boldsymbol{z}_{t,i}\|^{2}$ (23)
[…]
Cosine Similarity Loss: To further align the structure of features between the source and target domains, we treat the features as vectors in a high-dimensional space. A cosine similarity loss is proposed to achieve angular alignment between the source domain vectors and the target domain vectors.
$\mathcal{L}_{\cos}(\boldsymbol{z}_{i}^{S\leftrightarrow T})=\frac{\boldsymbol{z}_{s,i}\cdot\boldsymbol{z}_{t,i}}{\|\boldsymbol{z}_{s,i}\|\cdot\|\boldsymbol{z}_{t,i}\|}$ (24)
where $S$ and $T$ represent the source and target domains, respectively. Finally, the overall loss function can be described as follows:
$\mathcal{L}_{\text{total}}(\boldsymbol{Z}^{S},\boldsymbol{Z}^{T})=\alpha\mathcal{L}_{\text{cos}}(\boldsymbol{Z}^{S},\boldsymbol{Z}^{T})+\beta\mathcal{L}_{\text{mse}}(\boldsymbol{Z}^{S},\boldsymbol{Z}^{T})$ (25)
where $\boldsymbol{Z}^{S}$ and $\boldsymbol{Z}^{T}$ represent all feature vectors extracted from the source domain and target domain, respectively, while $\alpha$ and $\beta$ are hyper-parameters controlling the effect of the two loss functions.
Building upon the loss function of the transfer learning approach, the optimization of the target domain model parameters $\boldsymbol{\psi}_{\text{t}}$ is performed using first-order methods, e.g., Adaptive Moment Estimation (Adam) optimizer. To provide a more precise depiction of the overall process, the transfer learning procedure of DOA-ViT using Adam is detailed as Algorithm 1.
Algorithm 1 Transfer learning Approach using Adam
0:  DOA-ViT weights $\boldsymbol{\psi}_{s}$ pretrained on the source domain, hyperparameter $\epsilon$.
0:  Target domain Dataset $\mathcal{D}_{\text{target}}$, learning rate $\mu$, number of batches $B$, epochs number $e_{\max}$.
1:  Initialize DOA-ViT weights $\boldsymbol{\psi}_{s}$ for source domain model.
2:  Initialize DOA-ViT weights $\boldsymbol{\psi}_{t}$ for target domain model.
3:  for $\text{epoch}=0,1,\dots,e_{\max}$ do
4:     Randomly divide $\mathcal{D}$ into $B$ batches $\{\mathcal{D}_{b}\}_{b=1}^{B}$.
5:     for $b=1,2,\dots,B$ do
6:        for $(\widetilde{\boldsymbol{R}}_{b}^{(j)},\boldsymbol{\theta}_{b}^{(j)})\in\mathcal{D}_{b}$ do
7:           A Simulation Data Generator creates ideal covariance matrices $\widetilde{\boldsymbol{R}}_{b}^{*}$ based on $\boldsymbol{\theta}_{b}^{(j)}$
8:           Compute representations $\boldsymbol{Z}^{S}(\widetilde{\boldsymbol{R}}_{b}^{(j)},\boldsymbol{\psi}_{s})$ for the source domain.
9:           Compute representations $\boldsymbol{Z}^{T}(\widetilde{\boldsymbol{R}}_{b}^{*},\boldsymbol{\psi}_{t})$ for the target domain.
10:        end for
11:        Compute $\mathcal{L}_{\mathcal{D}}(\boldsymbol{Z}^{S},\boldsymbol{Z}^{T})$ through Equation 25.
12:        Update weights via $\boldsymbol{\psi}_{t}\leftarrow\boldsymbol{\psi}_{t}-\mu\nabla_{\psi}\mathcal{L}_{\mathcal{D}_{b}}(\boldsymbol{Z}^{S},\boldsymbol{Z}^{T})$.
13:     end for
14:  end for
15:  return  $\boldsymbol{\psi}_{t}$
V SIMULATIONS AND ANALYSES
As the DOA estimation system is a typical electronic system, the performance of the model in DOA estimation is influenced by a combination of multiple parameters. We summarize the primary factors as follows:
  1. AS1
    SNR and the number of snapshots. These two parameters are the most extensively discussed in DOA estimation fields. It is generally accepted that high SNR and a large snapshots lead to an ideal signal environment, thereby enhancing algorithm performance.
  2. AS2
    Number of array elements $\boldsymbol{M}$ and number of incident signals $\boldsymbol{K}$. A larger $M$ and a smaller $K$ facilitate more accurate DOA estimation.
  3. AS3
    Angular Separation and Range of incident signals. Discussions regarding the angular parameters of incident signals are relatively limited; however, these factors significantly impact a model’s performance. In general, smaller angular separations result in more similar steering vectors, leading to increasingly difficult DOA estimation. This situation poses a considerable challenge to the model’s angular resolution. Additionally, the range of incoming signals can affect the DOA estimation performance of specific array types (such as $\theta$ for ULA or $\phi$ for UCA); greater angles of signal incidence lead to reduced angular resolution for these arrays.
In this section, multiple distinct simulation scenarios were constructed based on the settings of the aforementioned parameters. We demonstrate the effectiveness of the proposed model and transfer learning approach through extensive simulation experiments. We begin with a detailed description of the experimental setup in Subsection V-A. Subsequently, the performance of our proposed model under these setups is reported in Subsection V-B. Furthermore, we demonstrate that our proposed model exhibits significant performance degradation in scenarios involving array imperfections, which motivates the introduction of transfer learning techniques to address the challenges posed by array imperfections. The simulation results are presented in Subsection V-C.
A wide range of classic and state-of-the-art algorithms were selected to demonstrate the effectiveness of the proposed model. We first selected several classic subspace-based approaches as baseline methods for comparison, such as MUSIC [17] and Unity-ESPRIT algorithms [16]. Compressive sensing algorithm $\ell_{1}$-SVD algorithm [24, 63] is considerd, which has strong performance in low SNR and small snapshots scenarios. In addition, a series of deep learning methods and their variants that integrate traditional algorithms were used for comparative analysis. We selected spatial spectrum estimation CNN algorithm (SPE-CNN) [1], SubspaceNet [35], ASL-2 [33], and Learning-SPICE [34] for comparison. SPE-CNN is recognized as a well-established deep learning-based method and SubspaceNet can be understood as a subspace algorithm implemented using deep learning networks. Furthermore, ASL-2 and Learning-SPICE, which representing the latest algorithms that combine deep learning and compressive sensing, are expected to perform well in small snapshots scenarios.
V-A Experimental Setup
1) Simulation Scenarios: Throughout this study, we consider four distinct scenarios to evaluate the performance of our proposed algorithm:
  • Scenario 1: A ULA with $M=8$ and $K=3$.
  • Scenario 2: A ULA with $M=8$ and $K=7$.
  • Scenario 3: A ULA with $M=16$ and $K=3$.
  • Scenario 4: A UCA with $M=12$ and $K=5$.
Figure 4: DOA estimation under Scenarios 1 with $SNR=-5db$ and $snap=10$. Three random incident signals are generate and we focus on the first one. The points represent the estimated DOA provided by the algorithms while the circles denote the true DOAs.
TABLE I: CDF-Based Statistical Evaluation results under Scenarios 1 with $snap=10$ and $SNR=-5$.
Algorithms Ours MUSIC Unity-ESPRIT $\ell_{1}$-SVD SPE-CNN ASL-2 SubspaceNet Learning-SPICE
$\text{CDF}(3)$ 55.36% 54.10% 48.90% 44.30% 60.11% 33.49% 52.63% 68.08%
$\text{CDF}(10)$ 89.20% 73.43% 80.31% 82.93% 81.35% 55.68% 87.35% 87.51%
CDF10% $0.455^{\circ}$ $0.410^{\circ}$ $0.526^{\circ}$ $0.638^{\circ}$ $0.368^{\circ}$ $0.766^{\circ}$ $0.501^{\circ}$ $\mathbf{0.309^{\circ}}$
CDF90% $\mathbf{10.63^{\circ}}$ $38.13^{\circ}$ $24.22^{\circ}$ $20.91^{\circ}$ $29.66^{\circ}$ $55.20^{\circ}$ $13.55^{\circ}$ $16.94^{\circ}$
Scenario 1 represents a standard multi-target DOA estimation scenario, which is widely adopted for evaluating model performance. Subsequently, to further investigate whether the model’s performance deteriorates when handling a larger number of incident angles, Scenario 2 is introduced.
Scenario 3 represents a scenario with more array elements, which aims to achieve precise DOA estimation. In such cases, algorithms typically demand significantly greater computational resources. Thus, maintaining accuracy while minimizing computational overhead becomes a critical challenge in this scenario.
Scenario 4 involves the use of a UCA for two-dimensional localization. This scenario evaluates the model’s capability to extend from one-dimensional to two-dimensional tasks, and tests its performance in 2D DOA estimation.
The aforementioned scenarios essentially modify the parameters of AS2. Within each scenario, data with varying SNR, snapshot numbers, angular separations and angular ranges are thoroughly tested. This comprehensive evaluation aims to provide a thorough assessment of the model’s performance, including model’s adaptability to low SNR and small snapshots scenarios, model’s angular resolution capabilities, and model’s angular estimation range in various scenarios.
2) Hyperparameter Settings: In the developed DOA-ViT model, the dimensionality of the embedded vectors $D$ is fixed at 768, while the dimension of each head $d_{k}$ is set to 64. Additionally, our model consists of 6 Transformer encoders and the MHSA module is configured with 12 attention heads. Within each encoder, the expansion factor of the MLP layers is set to 4.
Under each configuration, 50,000 samples are generated for model training, and 20,000 samples are generated for model validation. For a ULA, the incident angles $\theta$ are randomly generated within the range $[-60^{\circ},60^{\circ}]$ in the training and validation sets. And for a UCA, the incident angles $\theta$ are randomly generated within $[-180^{\circ},180^{\circ}]$, and $\phi$ values are randomly generated within $[0^{\circ},60^{\circ}]$.
The Adam optimizer is configured with a learning rate of $0.0001$; betas values $(0.9,0.999)$ are used for the first and second moment estimates, which is considered effective in most scenarios. The number of training epochs is set to 500. However, if the validation loss does not improve over the course of 30 epochs, early stopping is triggered to halt the training.
For the sake of simplicity and fairness, we train each deep learning model for comparison using the same dataset and optimizer to ensure consistency in the experiments and comparability of the results. For more detailed implementation and configuration specifics, the complete code is available in our open-source repository [^1] https://github.com/zzb-nice/DOA_est_Master, providing further insight into the structure and parameter settings employed.
V-B DOA estimation performance based on ViT model
V-B1 Simulation results in Scenario 1
A series of simulation experiments were conducted to validate the effectiveness of the proposed ViT model. First, we evaluate the performance of the model in Scenario 1.
Visual Performance Demonstration: In order to provide a visual comparison of the performance of various algorithms in Scenario 1, specifically in the low SNR and small snapshot conditions, we visually demonstrate the DOA estimation results of multiple algorithms in Fig 4. Scenario 1 with SNR set to -5 dB and snapshots number set to 10 is considered, which serves as the low SNR, small snapshot scenario. The absolute angular error of the model’s DOA estimates is plotted on the right side. Through a direct comparison, the proposed method demonstrates superior performance, with the majority of estimation errors being smaller than $25^{\circ}$ in such a low SNR, small snapshot scenario. In contrast, all other algorithms exhibit severe deviations with absolute errors exceeding 50 degrees.
Figure 5: RMSE $vs.$ SNR results under Scenarios 1 with $snap=10$. The incident signal’s DOA values are randomly generated with the minimum angular interval set to 3 degrees.
RMSE $\boldsymbol{vs.}$ SNR: To effectively evaluate the model’s performance, it is essential to assess its statistical performance. We demonstrate the Root Mean Squared Error (RMSE) of estimated DOA values under varying SNR and snapshots (AS1).
In Fig. 5, we investigate the RMSE vs. SNR results based on 5,000 simulations. It can be seen that the proposed method achieves the best performance when the incident angles are randomly generated, which can be regarded as the most general case. Learning-SPICE closely approaches the performance of the proposed model as SNR increase ($SNR=0db$ or $SNR=5db$). Unity-ESPRIT and SPE-CNN perform relatively well at high SNRs, while SubspaceNet achieves better results in low SNR scenarios. $\ell_{1}$-SVD exhibits consistently stable performance, and MUSIC and ASL-2 exhibit inferior performance compared with the aforementioned algorithms.
Furthermore, in order to provide a more intuitive demonstration of the model’s statistical performance and a more comprehensive evaluation of its overall performance, the cumulative distribution function (CDF) of the DOA estimation absolute error is illustrated to assess the model’s performance, which describes the probability distribution of the absolute error. As shown in Fig. 6, 5,000 simulations are conducted in each condition, where the incident signal’s DOA values are randomly generated with the minimum angular interval set to 3 degrees.
Through the CDF, it can be observed that under higher SNR conditions (e.g., $SNR=0db$ or $5db$), the proposed model, Unity-ESPRIT, SPE-CNN and Learning-SPICE exhibited commendable performance. However, compared with other algorithms, the proposed model exhibits a lower cumulative probability of small absolute errors, and it shows a distinct advantage in larger absolute error ranges. This trend is more evident in low SNR scenarios like $SNR\leq 0db$, as detailed in the following analysis.
(a) $SNR=0db$
(b) $SNR=5db$
(c) $SNR=-5db$
(d) $SNR=-10db$
Figure 6: CDF of the absolute error under Scenarios 1 with $snap=10$ and various $SNR$, for (a) $SNR=0db$, (b) $SNR=5db$, (c) $SNR=-5db$ and (d) $SNR=-10db$.
A representative scenario is depicted in Fig. 6, where it can be observed that the distribution of absolute errors within the range of $0^{\circ}$ to $3^{\circ}$ for the proposed model is slightly lower than those of Learning-SPICE, SPE-CNN and MUSIC, while exhibiting advantage over the SubspaceNet, Unity-ESPRIT, $\ell_{1}$-SVD and ASL-2 algorithms. However, the proposed model’s probability of absolute errors exceeding $10^{\circ}$ is substantially lower, which indicates a reduced frequency of high-error occurrences, highlighting our model’s superior robustness relative to other approaches.
Figure 7: RMSE $vs.$ SNR results under Scenarios 1 with $snap=10$. The dashed line in the figure represents Scene 1, while the solid line represents Scene 2. In Scene 1, the true DOA values are [10°, 13°, 16°], and in Scene 2, the true DOA values are [10°, 20°, 30°].
Figure 8: RMSE $vs.$ SNR results under Scenarios 2 with $snap=10$. The incident signals are randomly generated with the minimum angular interval set to 3 degrees.
Figure 9: RMSE $vs.$ snap results under Scenarios 1 with $SNR=-10db$. The dashed line in the figure represents Scene 1, while the solid line represents Scene 2. In Scene 1, the true DOA values are [5.1°, 30.3°, 54.5°], and in Scene 2, the true DOA values are [-9.9°, 15.3°, 39.5°].
Figure 10: RMSE $vs.$ SNR results under Scenarios 3 with $snap=10$. The incident signals are randomly generated with the minimum angular interval set to 3 degrees.
Furthermore, through the CDF, a series of statistical measures related to the absolute error can be introduced. For example, CDF50% represents the absolute error corresponding to a cumulative probability of ${50\%}$. As a quantile-based measure of absolute error, it differs from the commonly used mean absolute error (MSE), which represents the mean of absolute errors. Both serve as statistical measures of absolute error and hold statistical significance. Similarly, CDF(5) refers to the probability that the absolute error is less than 5, which can also reflect the model’s performance.
The results of the CDF-based statistical evaluation for various models are presented in Table I. From the table, it can be observed that Learning-SPICE demonstrates superior performance at $\text{CDF}(3)$ and CDF10%, achieving the highest DOA estimation accuracy, where the algorithm’s estimated values are most likely to be close to the true values. However, our proposed model demonstrates the best performance at $\text{CDF}(10)$ and CDF90%, showing greater robustness by reducing the occurrence of large errors. Due to the fact that RMSE penalizes larger estimation errors severely, the proposed model demonstrates superior performance in terms of RMSE compared with the Learning-SPICE algorithm.
To further demonstrate the performance of the proposed model, we selected four representative models mentioned above for testing in smaller angular interval scenarios. We consider two Scenes with different angular separations (AS3): Scene 1 represents the scenario with signals incident at [10°, 13°, 16°], while Scene 2 represents the scenario with signals incident at [10°, 20°, 30°]. In Scene 1, the angular interval between incident signals is smaller, which imposes higher demands on the model’s angular super-resolution capability.
In general, since smaller angular intervals result in more similar steer vectors, algorithms based on linear space and compress sensing will perform better in Scene 2 than in Scene 1. On the other hand, deep learning algorithms are fundamentally designed to minimize the loss function. By minimizing the loss function across the entire training set, deep learning algorithms may achieve improved performance at smaller angular intervals.
[…]
RMSE $\boldsymbol{vs.}$ snapshots: To further demonstrate the model’s performance under varying snapshots (AS1) and angular ranges (AS3) for a ULA, we conducted simulation experiments in different snapshot scenarios. Similar to the previous experiment, we empirically evaluate the RMSE by averaging over 5,000 Monte Carlo trials. Two distinct scenarios are considered to investigate the model’s performance under different angular ranges. Scene 1 represents the scenario with signals incident at [5.1°, 30.3°, 54.5°], while Scene 2 represents the scenario with signals incident at [-9.9°, 15.3°, 39.5°]. The demands on the model’s angular super-resolution are alleviated and the angular intervals are kept the same in two scenes. Thus, the larger angular range in Scene 2 leads to a lower angular resolution, which has a negative impact on the performance of the algorithms.
As shown in Fig. 10, a clear trend can be observed: for any given algorithm, the performance in Scene 1 is inferior to that in Scene 2, which can be attributed to the more challenging conditions in Scene 1. Furthermore, the proposed DOA-ViT model generally outperforms the other algorithms in this low SNR scenario. In Scene 1, our model achieves the best performance, surpassing the results achieved by MUSIC, $\ell_{1}$-SVD, Learning-SPICE, and Unity-ESPRIT. In Scene 2, the proposed model also attains the best performance, slightly outperforming the Learning-SPICE algorithm. These results highlight the superior effectiveness of our approach.
V-B2 Simulation results in Scenario 2 and 3
(a) ViT under Scene 1
(b) MUSIC under Scene 1
(c) ViT under Scene 2
(d) MUSIC under Scene 2
Figure 11: Scatter plots of our proposed model and MUSIC algorithm under Scenarios 4 with $snap=50$ and $SNR=-5db$. Scene 1 is configured with $\boldsymbol{\theta}=[-120,-60,0,60,120]$ and the corresponding azimuth angles $\boldsymbol{\phi}=[30,10,30,10,35]$. Scene 2 is configured with $\boldsymbol{\theta}=[-135,-120,-60,0,60]$, $\boldsymbol{\phi}=[35,30,10,30,10]$.
(a) CDF of $\boldsymbol{\theta}$
(b) CDF of $\boldsymbol{\phi}$
Figure 12: The CDF of the absolute error under Scenarios 4 with $snap=50$ and $SNR=-5db$ is shown for: (a) $\boldsymbol{\theta}$, and (b) $\boldsymbol{\phi}$. The angles of the incident signals are randomly generated.
To evaluate the model’s performance in handling a larger number of incident angles or with more array elements, we extend the testing from Scenario 1 to Scenario 2 and Scenario 3.
We plotted the RMSE $vs.$ SNR results under Scenario 2 with $snap=10$ in Fig. 10. In this case, the incident signals are randomly generated within the range of $[-60^{\circ},60^{\circ}]$, and the minimum angular interval is set to 3 degrees. If the generated angular interval is less than 3 degrees, the incident angle will be regenerated. Notably, it can be observed that, compared with all the subspace-based methods, compressive sensing methods, and the state of the art deep learning algorithms, the proposed approach achieves superior statistical performance in low SNR and small snapshots scenarios. This indicates that the proposed model can still deliver relatively robust performance even in scenarios involving a greater number of incident angles.
Furthermore, the algorithm is tested under Scenario 3 with $\text{snap}=10$. In Fig. 10, it can be observed that the proposed model outperforms other algorithms in low SNR and small snapshot scenarios, except for cases where $SNR=-5db$ or $SNR=5db$. $\ell_{1}$-SVD and Unity-ESPRIT algorithms demonstrates high accuracy in scenarios with a larger number of array elements. Due to the limitations of data-driven algorithm fitting accuracy, the proposed model could not show significant improvement in algorithm performance in the range of 0 dB to 5 dB. As a result, its performance is inferior to Unity-ESPRIT and $\ell_{1}$-SVD algorithms under the $SNR=5db$ scenario. The performance of the proposed model in DOA estimation is difficult to improve in high-precision and higher SNR scenarios, which is an area for future improvement.
V-B3 Simulation results in Scenario 4
To evaluate the model’s ability to extend from one-dimensional to two-dimensional scenarios, we expand the model to two dimensions by simultaneously regressing the values of $\theta$ and $\phi$, and model’s performance is tested in Scenario 4.
Similar to one-dimensional DOA estimation, two-dimensional DOA estimation is simultaneously influenced by multiple parameters. Therefore, we consider two Scenes with different angular separations (AS3). Since it is uncommon to simultaneously vary the angular separations of all signals in practice scenarios, we only modify one of the angular separation parameters across the two scenes. In Scene 1, the incident signals are generated with $\boldsymbol{\theta}=[-120,-60,0,60,120]$, with the corresponding azimuth angles given by $\boldsymbol{\phi}=[30,10,30,10,35]$. Scene 2 is configured with $\boldsymbol{\theta}=[-135,-120,-60,0,60]$ and $\boldsymbol{\phi}=[35,30,10,30,10]$.
A representative visualization of the results is presented in Fig. 11, which facilitates an intuitive comparison of the algorithm’s performance. In Fig. 11, the symbol ”+” represents the true location of the target, while the dot in the figure represents the predicted value of the algorithm. Additionally, the model’s predicted values and their corresponding true values are marked with the same color. 100 Monte Carlo simulations were conducted for each scene. According to the experimental results in Scene 1, it can be observed that the predicted results of the proposed model are predominantly clustered around the true values. In contrast, the MUSIC algorithm achieves better DOA estimation performance at the more widely separated targets, like $\theta=[-120,120]$ and $\phi=[30,35]$, but it tends to confuse the other three points, resulting in poorer prediction performance. In Scene 2, we moved the fifth target from $\theta=[120]$ $\phi=[35]$ to $\theta=[-135]$ $\phi=[35]$, which is adjacent to the first target. It can be observed that the MUSIC algorithm fails to distinguish between the two targets, forming a single spectral peak between them, whereas the predicted results of the proposed model remain scattered around the true values.
(a) $\rho=0$
(b) $\rho=0.5$
(c) $\rho=1$
Figure 13: Three different levels of array imperfections are proposed to evaluate the performance of our proposed model and traditional algorithms: (a) $\rho=0$, no array imperfections, (b) $\rho=0.5$, moderate array imperfections, and (c) $\rho=1$, severe array imperfections.
Figure 14: RMSE $vs.$ SNR results under Scenarios 1 with $snap=10$ and array imperfections of $\rho=1$. The transfer learning approach is employed to enhance the model’s performance under array imperfection scenarios. The dashed line represents the case where only 50 samples are used for transferring learning, while the solid line indicates the use of 500 samples.
Furthermore, 1000 simulations are performed under Scenario 4 with $snap=50$ and $SNR=-5\ \mathrm{dB}$ to evaluate the statistical performance of the proposed model, where the angles of the incident signals are randomly generated. The CDF of the proposed model is presented in Fig. 12. It can be observed that the MUSIC algorithm performs better on CDF(10). However, its performance deteriorates on CDF(30). This suggests that while the MUSIC algorithm is likely to achieve more accurate DOA estimations for certain targets, it also shows a higher probability of producing significant DOA estimation errors. In contrast, the proposed model demonstrates a more consistent overall stability in its predictions.
This experiment also demonstrates that the CDF of Absolute Error serves as an effective metric in DOA estimation tasks. Compared with RMSE, it provides a more comprehensive evaluation of the model’s performance. Supported by the Cramér-Rao bound as a theoretical benchmark, RMSE offers a significant reference value, whereas the CDF of Absolute Error achieves a more thorough assessment of the model’s capabilities. Through the simulation results, it is evident that the proposed model achieves a greater advantage in terms of RMSE. However, through the analysis of the CDF of Absolute Error, it has certain limitations results in the precision DOA estimation.
V-C Transfer Learning Approach for Array imperfection
The presence of array imperfections introduces inconsistencies between the received signal model and the ideal model, leading to a degradation in the DOA estimation performance of the model. Scenarios involving multiple array imperfections, particularly under low SNR and small snapshot conditions, are inherently complex and challenging to handle. In this section, we demonstrate the effectiveness of the proposed transfer learning approach in these scenarios through simulation experiments.
V-C1 Transfer Learning in Scenario 1
Figure 15: DOA estimation under Scenarios 1 with $SNR=5db$ and $snap=10$, $\rho=1$. Three incident signals are generated and the interval is set to $\Delta\hat{\theta}={10}^{\circ}$. The line represents the true location of the target, while the dot in the figure represents the predicted value of the algorithm. Additionally, the model’s predicted values and their corresponding true values are marked with the same color.
TABLE II: Transfer Learning Results with Varying Sample Number
Sample Number 20 50 100 200 300 500 800 1000
Directly train 23.03 23.44 23.04 23.03 23.00 14.43 23.00 22.87
Fine-tune 14.17 14.38 13.57 13.84 13.30 12.79 13.31 12.38
Trnasfer Learning 13.66 13.27 13.06 12.72 12.60 12.54 12.21 11.48
(a) ViT without transfer learning
(b) ViT with transfer learning
Figure 16: DOA estimation under Scenarios 1 with $SNR=5db$, $snap=10$ and array imperfections of $\rho=1$. 100 Monte-Carlo simulations are conducted and the incident signal is configured with $\boldsymbol{\theta}=[-135,-120,-60,0,60]$, $\boldsymbol{\phi}=[35,30,10,30,10]$.
To demonstrate the performance degradation of the deep learning model under array imperfections, we illustrate the model’s performance under scenarios with multiple array imperfections, as depicted in Fig. 13. The position bias, amplitude and phase inconsistency, and mutual coupling effects of the array systems are simultaneously considered, and the intensity of these array imperfections is controlled by a hyperparameter $\rho$. As shown in Fig. 13, It can be observed that under the influence of multiple array imperfections, the statistical performance of deep learning-based methods (the proposed model, Learning-SPICE and SPE-CNN), compressed sensing-based methods ($\ell_{1}$-SVD), and subspace-based methods (Unity-ESPRIT and MUSIC) all deteriorates noticeably. Although the proposed model achieves optimal performance under the ideal signal model, particularly in low SNR and small snapshots scenarios, its performance significantly degrades when subjected to array imperfections. This degradation becomes particularly pronounced in scenarios with relatively higher SNR. When $\rho=1$, the performance of the proposed model declines level similar to the traditional algorithms at $\rho=0$.
To validate the effectiveness of the transfer learning approach, we evaluate the performance of the model adapted through transfer learning under varying SNR through 5,000 Monte Carlo simulations. As shown in Fig. 14, It can be observed that the proposed model achieves significant performance even with only 50 samples, demonstrating accurate DOA estimation under low SNR and small snapshot scenarios. Moreover, when more samples (500 samples) are used, the model’s performance is further enhanced.
To demonstrate the superiority of the proposed transfer learning algorithm, a comparison experiment is carried out under Scenario 1 with $snap=10$, $SNR=-5$, and array imperfections $\rho=1$. Our method is compared with two other approaches: (1) directly training the model through the samples with array imperfections (Directly Train), and (2) fine-tuning the model trained on simulated data using samples with array imperfections (Fine-tune). The performance of these three methods are evaluated across varying numbers of array imperfection samples, as shown in Table II.
The results indicate that the proposed transfer learning algorithm consistently achieves superior performance in most cases. This demonstrates that a well-designed transfer learning strategy can utilize the information of limited samples more effectively, outperforming both the Fine-tune and Directly Train methods.
To illustrate the performance of the proposed model after transfer learning, an experiment is conducted to visually demonstrate our model’s performance, and the results are presented in Fig. 15. In this experiment, three incident angles are generated with a constant interval of $\Delta\hat{\theta}={10}^{\circ}$, covering the range from -60 to 60 degrees, with the incident angles linearly scaled with a step of 0.5 degrees. From the experimental results, it can be observed that the proposed model exhibits significant bias before applying transfer learning, particularly when the incident angles are close to -60 degrees. However, after applying transfer learning, the predict results of the proposed model become concentrated around the true values, demonstrating high DOA estimation accuracy. In contrast, due to the combined effects of multiple array imperfections, the performance of the MUSIC and $\ell_{1}$-SVD algorithm deteriorates, exhibiting substantial deviations. Meanwhile, Unity-ESPRIT and SPE-CNN show significant deviations when the incident angles span a wider range. SubspaceNet exhibits noticeable distortions, failing to achieve accurate DOA estimation. Learning-SPICE, SPE-CNN and ASL-2 demonstrate enhanced resilience to array imperfections; however, systematic distortion remains observable in the predict results.
V-C2 Transfer Learning in Scenario 4
To further validate the effectiveness of the proposed transfer learning algorithm, we extend the algorithm from one-dimensional to two-dimensional. The 2D DOA estimation results of the ViT before and after applying transfer learning are shown in Fig. 16. It can be clearly observed that prior to the application of the transfer learning algorithm, the model fails to effectively distinguish between the first target ($\theta=[-135]$, $\phi=[35]$) and the second target ($\theta=[-120]$, $\phi=[30]$) and exhibits significant bias for orther targets due to array imperfections. However, after the transfer learning algorithm is applied, the prediction bias was substantially corrected, and the predicted values were able to converge near the true target.
VI CONCLUSION
In this work, we develop a DOA-ViT model for DOA estimation, which leverages the attention mechanism to capture global dependencies. By directly learning the mapping from the sampled SCM to DOA values, the proposed method can be easily extended to arbitrary array geometries or 2D DOA estimation scenarios, demonstrating strong adaptability. Extensive simulations are conducted to evaluate the performance and robustness of the proposed model under various conditions: i) different SNRs and numbers of snapshots, ii) varying numbers of array elements $M$ and incident signals $K$, and iii) different angular separations and signal ranges.
Furthermore, deep learning based DOA estimators often suffer from severe performance degradation in practical scenarios with array imperfections. To overcome this limitation, we propose a supervised transfer learning strategy to mitigate the performance degradation in practical scenarios. By aligning the features of the source and target domains in a supervised manner, the proposed method demonstrates improved results in scenarios with array imperfections.
It is worth noticing that this work is the first to introduce an advanced transfer learning strategy to address array imperfections, rather than simply employing fine-tune. Investigating how to effectively leverage transfer learning methods to improve the performance of deep learning models in scenarios with array imperfections is a significant and worthwhile research direction, and has the potential to enhance the performance of deep learning algorithms in practical applications. Further exploration of various transfer learning approaches for handling array imperfections is encouraged.
ACKNOWLEDGMENT
The authors would like to thank the anonymous reviewers for their insightful comments and suggestions.

### v5

@@
-A Deep Learning-Based Supervised Transfer Learning Framework for DOA Estimation with Array Imperfections
+TransDOA: Calibrating Array Imperfections via Transformer-based Transfer Learning
 Abstract
@@
 In the proposed approach, we highlight three major contributions.
-First, we propose a Vision Transformer (ViT) based method for DOA estimation, which achieves excellent performance in scenarios with low signal-to-noise ratios (SNR) and limited snapshots.
+First, we propose a Transformer based method for DOA estimation, which achieves excellent performance in scenarios with low signal-to-noise ratios (SNR) and limited snapshots.
 Second, we introduce a transfer learning framework that extends deep learning models from ideal simulation scenarios to complex real-world scenarios with array imperfections.
@@
 Our code can be accessed at https://github.com/zzb-nice/DOA_est_Master.
-Index Terms:
-Direction of Arrival (DOA) estimation, Domain Adaptive, Vision Transformer, Data-driven Methods, Deep learning
 I INTRODUCTION
-Direction of Arrival (DOA) estimation finds extensive applications across various engineering fields, including wireless communications, astronomical observation, radar, and sonar [2, 3, 4, 5, 6], where accurately determining the direction of incoming signals is crucial for efficient operation.
-Today, as radar technology becomes increasingly accessible for civilian use, DOA estimation algorithms have also been applied in emerging areas such as autonomous driving and drone operating system [7, 8, 9, 10].
-With the continuous development of engineering technologies, practical scenarios are placing higher demands on DOA estimation algorithms.
-For example, in autonomous driving scenarios, there is an increasing need for DOA estimation of fast-moving small targets [7, 11].
-In non-cooperative target localization scenarios, it is essential to achieve rapid and precise DOA estimation [12, 13].
-These practical scenarios generally require accurate DOA estimation results under conditions of low SNR and small snapshots, along with high real-time performance.
-Therefore, there is an urgent need for a new algorithm with relatively low computational complexity, one that can achieve strong performance in ideal scenarios while also demonstrate improved performance in low SNR and small snapshots scenarios.
-Over the past few decades, extensive research has been conducted on DOA estimation, leading to the development of numerous algorithms [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 1, 29, 30, 31, 32, 33, 34, 35].
-These methods can be broadly categorized into model-driven and data-driven approaches.
-Among them, model-driven methods have been more extensively studied due to their well-established theoretical foundations.
-The core principle of model-driven DOA estimation is to establish a mapping between the directions of incoming signals and the received array data, and then use a series of techniques to accurately invert this mapping based on the received data.
-Based on different estimation strategies, model-driven approaches can be further classified into several major categories, including beamforming methods [14], subspace methods [15, 16, 17, 18, 19], maximum likelihood methods [20, 21, 22], and compressed sensing (sparsity-inducing) methods [36, 23, 24, 25].
-A leading and well-respected family of model-driven approaches is the subspace method, which achieves satisfactory performance in scenarios with high signal-to-noise ratio (SNR) and large snapshots.
-It maintains relatively low computational complexity when a limited number of array elements is used.
-However, its performance deteriorates significantly in practical scenarios with low SNR and small snapshots.
-Moreover, subspace-based methods struggle with color noise, and require the estimated number of signal sources to be smaller than the number of array elements, which limits their applicability.
-Compressive sensing-based and maximum likelihood-based methods have achieved certain improvements but are limited by real-time performance constraints.
-In summary, model-driven methods for DOA estimation have reached a performance bottleneck in practical scenarios.
+Figure 1: Schematic illustration of environmental perception using front-mounted mmWave radar in a night-time autonomous driving scenario.
+The radar point clouds are projected onto the camera image for visualization.
+Direction of Arrival (DOA) estimation plays a pivotal role across diverse engineering disciplines, including wireless communications, radio astronomy, radar, and sonar [1, 2, 3, 4], where precise signal localization is critical.
+Recently, the proliferation of accessible wireless sensing technology has expanded the scope of DOA algorithms to emerging commercial domains, such as autonomous driving and unmanned aerial systems (UAS) [5, 6, 7, 8].
+Rapid advancements in engineering have imposed increasingly stringent requirements on DOA estimation algorithms.
+For instance, autonomous driving necessitates the tracking of fast-moving targets with small radar cross-sections [5], while non-cooperative target localization demands both high speed and precision [9].
+In practice, these tasks are significantly complicated by the coexistence of challenging signal environments (e.g., low SNR and limited snapshots) and inevitable hardware impairments.
+Fig.
+1 illustrates the impact of signal-level and hardware-level challenges [10].
+In this real-world autonomous driving scenario, their interplay leads to severe performance degradation, manifested as significant point cloud sparsity and missed detections (e.g., the sparsely represented vehicle in the red box).
+However, existing algorithms often fail to address these coupled constraints simultaneously.
+Therefore, there is a compelling need for a new algorithm that delivers robust performance against array mismatches while maintaining high accuracy in low SNR and limited snapshot regimes.
+Existing DOA estimation methods can be broadly categorized into model-driven and data-driven approaches.
+Among them, model-driven methods have been more extensively investigated due to their well-established theoretical foundations.
+Based on different estimation strategies, model-driven approaches can be further classified into several primary categories, including beamforming methods [11], subspace-based methods [12, 13, 14], maximum likelihood methods [15, 16], and compressed sensing (sparsity-inducing) methods [20, 17, 18, 19].
+A leading and well-respected family of model-driven approaches is the subspace-based methods, which achieve satisfactory performance in scenarios with high SNR and large snapshots.
+However, their performance relies heavily on the asymptotic accuracy of the covariance matrix, which drastically degrades under low SNR and limited snapshot conditions.
+Moreover, these methods struggle with color noise, and require the estimated number of signal sources to be smaller than the number of array elements, which limits their applicability.
+Even if these signal-level limitations could be overcome, a more fundamental, physical challenge remains: the divergence between the idealized model and the physical hardware.
+Most standard model-driven methods typically rely on an ideal array manifold.
+However, The existence of array position biases, amplitude and phase inconsistencies, mutual coupling [21], and other array imperfections can severely distort the received data.
+In such cases, the mismatch between the theoretical model and the physical array often leads to significant performance degradation or even the complete failure of DOA estimation.
+To mitigate these hardware impairments, various active calibration [22, 23, 24] and self-calibration [25, 26, 27, 28, 29, 30, 31, 32] techniques have been developed.
+Active calibration methods, while affording high accuracy through the utilization of predefined auxiliary sources, necessitate significant resource expenditure and introduce operational discontinuity.
+In contrast, self-calibration methods offer a more practical solution by estimating array parameters and DOAs simultaneously without auxiliary sources.
+However, these methods often rely heavily on initialization and struggle to handle more severe array imperfections.
+As noted in [25], most existing self-calibration methods are ineffective in practical scenarios with low SNR, limited snapshots, and spatially proximate sources——precisely the challenges that many DOA estimation algorithms face.
+Moreover, existing calibration methods often lack versatility.
+Due to the complexity of modeling combined imperfections, most algorithms focus on correcting only a single type of array imperfections, such as array position biases [22, 24], amplitude-phase inconsistencies [26, 27, 28, 29], or mutual coupling [30, 31, 32].
+Only a few studies have explored to address multiple simultaneous imperfections [25, 23], and these often rely on specific array geometries (e.g., ULA/UCA) or other restrictive assumptions [25].
+Consequently, developing a robust framework capable of handling multiple array imperfections simultaneously, especially under low SNR and limited snapshot regimes, remains an urgent and open research problem.
 In recent years, the rapid advancement of computational hardware [33] has significantly accelerated the development of deep learning technologies.
-As data-driven technologies, machine learning and deep learning methods have gained increasing attention in the field of DOA estimation.
-These methods build complex mapping relationships from input to output based on large-scale datasets.
-Unlike model-driven methods that rely on predefined signal models, data-driven methods generally do not impose prior assumptions on the input and output.
-This makes them particularly suitable for practical scenarios where the signal model may be incomplete or inaccurate.
-Furthermore, during practical implementation, data-driven methods demonstrate better real-time performance due to their fast inference speed.
+Empirical studies have demonstrated the superior capability of deep learning-based methods to approach the Cramér-Rao Lower Bound (CRLB) in challenging signal regimes characterized by low SNR and limited snapshots [34, 35].
 As a result, data-driven methods are considered promising tools for achieving DOA estimation in practical scenarios.
-Deep learning methods were first introduced to address the problem of DOA estimation in the 1990s [26, 27].
-Following a considerable period of minimal utilization, these approaches have witnessed a substantial resurgence in DOA estimation research during recent years.
-The investigations conducted in [28, 1] pioneered the implementation of Deep Neural Network (DNN) and Convolutional Neural Network (CNN) for DOA estimation.
-Subsequently, the method proposed in [29, 30] advanced the field by first partitioning the incoming signal into multiple sub-regions through a network architecture, followed by the deployment of parallel networks across the sub-regions.
-Additionally, some methods aim to combine traditional DOA estimation techniques with deep learning methods.
-Due to their inherent flexibility, deep learning methods can effectively establish diverse mapping relationships.
-As a result, deep learning techniques are not necessarily limited to direct mapping to the DOA values but can also be applied to map various variables, such as noise-free covariance matrices, signal subspaces, angular intervals of incident signals, and so on.
-This flexibility makes them well-suited for integration with traditional methods such as subspace techniques and compressed sensing.
-In 2022, the method proposed in [31, 32] integrated subspace-based methods with deep neural networks.
+Additionally, recent studies have concentrated on combining traditional DOA estimation methodologies with deep learning-based architectures to realize further advancements in performance.
+In 2022, the method proposed in [36] integrated subspace-based methods with deep neural networks.
 Specifically, a CNN was employed to reconstruct a noise-free covariance matrix, after which a classical subspace-based approach was applied for DOA estimation.
@@
 Unlike previous work, SubspaceNet utilized the differentiability of polynomial equation root-finding and matrix decomposition, which allowed for the backpropagation of loss from the Root MUSIC algorithm.
-Nevertheless, most of the model-driven and data-driven methods mentioned above do not account for array imperfections in practical systems.
-The existence of array position biases, amplitude and phase inconsistencies, mutual coupling effects [38], and other array imperfections can affect the DOA estimation accuracy of models.
-In severe cases, these imperfections can significantly degrade the model’s DOA estimation performance or even cause its DOA estimation to fail.
-A series of active calibration [39, 40, 41, 42, 43] and self-calibration [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54] methods have been applied to compensate for array imperfections in practical scenarios.
-Among them, active calibration methods require the setup of known signal sources with predefined DOA directions, which are specifically employed for the compensation of array imperfections.
-In contrast, self-calibration methods do not require additional auxiliary signal sources and can dynamically calibrate the array imperfections during array system applications, offering significant practical advantages.
-However, active calibration methods benefit from utilizing known DOA values for array imperfection estimation, leading to higher accuracy.
-In contrast, self-calibration methods generally estimate array imperfections simultaneously with DOA values, often relying on initialization and struggling to handle more severe array imperfections.
-As noted in [44], most existing self-calibration methods can perform well only under relatively ideal signal conditions and are ineffective in practical scenarios with low SNR, small snapshots, and spatially proximate sources——precisely the challenges that many DOA estimation algorithms face.
-Moreover, existing calibration methods are constrained by numerous limitations.
-Due to the significant difficulty in directly modeling and analyzing all array imperfections, most existing calibration algorithms can focus on only one of the array imperfections, such as array position biases [39, 42, 43], amplitude-phase inconsistencies [45, 46, 47, 48, 49] or mutual coupling [50, 51, 52, 53, 54].
-Only a few studies have explored how to address scenarios involving multiple array imperfections [44, 40, 41].
-Some methods are even restricted to calibrating specific array structures such as uniform linearity/circularity array [54, 48, 49, 53], or impose additional assumptions to simplify the model [44].
-Therefore, how to effectively address scenarios with multiple array imperfections remains an issue that requires further exploration.
-In [29], Liu et al.
-mentioned that deep learning algorithms exhibit good adaptability to array imperfections, but this adaptability fails in practical scenarios with low SNR and small snapshots.
-Based on the simulation results, it can be observed that in scenarios with low SNR and small snapshots, deep learning models remain sensitive to array imperfections, requiring the application of calibration models to compensate for array imperfections.
-However, due to the inherent flexibility of deep learning, it may be possible to directly construct a transfer learning algorithm [55] that transitions from an ideal scenario without array imperfections to a practical scenario with array imperfections, thereby enhancing the model’s adaptability to practical scenarios.
+Although these deep learning architectures have proven highly effective in challenging signal environments, the potential of utilizing deep learning techniques to mitigate physical array imperfections remains significantly unexplored in existing literature.
+While Liu et al.
+[35] suggested that deep learning algorithms exhibit good adaptability to array imperfections, this adaptability does not extend to all operating conditions.
+As will be empirically demonstrated in Fig.
+12 and Fig.
+14, deep learning models exhibit a pronounced sensitivity to array imperfections in challenging signal environments.
+Therefore, developing a robust deep learning-based approach capable of handling the simultaneous coexistence of adverse signal conditions and inevitable hardware impairments remains a critical and unresolved research challenge.
+As an emerging deep learning technique, transfer learning algorithm [40] is capable of leveraging knowledge acquired from a source task to enhance the performance of a distinct target task.
+This characteristic holds the potential to enhance algorithmic resilience in practical scenarios, particularly where challenging signal environments and inevitable hardware impairments coexist.
 Nonetheless, there is currently a lack of research on the application of transfer learning methods in the field of DOA estimation.
@@
 applied transfer learning to enhance the network’s adaptability to array imperfections, but they only fine-tuned the network from ideal data to array imperfections-affected data, without incorporating more advanced transfer learning techniques.
-In this paper, we propose a deep learning-based supervised transfer learning framework for DOA estimation, which aims to mitigate the performance degradation of data-driven approaches in practical scenarios with various array imperfections.
+In this paper, we propose a Transformer-based transfer learning framework for DOA estimation, which aims to mitigate the performance degradation of data-driven approaches in practical scenarios with various array imperfections.
 In our approach, the data is divided into source and target domains.
 The source domain data consists of ideal data without array imperfections, while the target domain data represents practical scenarios containing array imperfections.
-Firstly, we train a DOA estimation model using the source domain data.
-A transfer learning method is then employed to align the features between the source and target domains, thereby enhancing the deep learning model’s ability to adapt to array imperfections and improving performance in scenarios with low SNR, small snapshots, and multiple array imperfections.
-Moreover, in a variety of scenarios, extensive experiments have been conducted to validate the effectiveness of the proposed model.
+Firstly, we propose a DOA estimation model, termed TransDOA, which is trained initially on the source domain data.
+A transfer learning method is then employed to align the features between the source and target domains, thereby significantly enhancing TransDOA’s ability to adapt to array imperfections and improving performance in scenarios with low SNR, limited snapshots, and multiple array imperfections.
+Moreover, in a variety of scenarios, extensive experiments have been conducted to validate the effectiveness of the proposed approach.
 The contributions of our work can be outlined as follows:
   1. 1.
-    We propose a novel DOA estimation model based on Vision Transformer, which demonstrates exceptional performance under challenging conditions, such as low SNR and small snapshot scenarios.
+    We propose a novel DOA estimation model, TransDOA. This model demonstrates exceptional performance under challenging conditions, such as low SNR and limited snapshot scenarios.
   2. 2.
     Due to the presence of array imperfections, the data distributions of the source and target domains differ significantly, leading to substantial performance degradation when the model is deployed in practical scenarios.
-To address this issue, we introduce a transfer learning algorithm to align the features between the source and target domains, enhancing the model’s performance in practical scenarios.
+To address this issue, we introduce a transfer learning algorithm to align the features between the source and target domains, enhancing TransDOA’s performance in practical scenarios.
   3. 3.
     Via extensive simulations, we compare the proposed method with existing approaches across multiple evaluation metrics and demonstrate the superiority of our method in terms of DOA estimation accuracy and robustness.
+This paper is organized as follows.
+Section II formulates the problem.
+Section III presents the design of a network for DOA estimation.
+Section IV introduces a transfer learning scheme for the developed model.
+Experimental results are given in Section V.
+Finally, Section VI concludes the paper.
 II Problem Formulation
-Figure 1: The framework of DOA-ViT and transfer learning approach proposed for array imperfections.
-In this work, two models of array receiving signals are considered.
-The first is the ideal array signal model, where the signal received by each sensor is represented as $Be^{j\varphi}s(t)$.
-where $B$ denotes the sensor’s gain for the incoming signal, and $e^{j\varphi}$ accounts for the phase shift caused by the time delay of the signal propagation between the sensors.
-This idealized model aligns with the standard assumptions used in most DOA estimation methods.
-The second signal model incorporates array imperfections to reflect real-world conditions.
-It considers factors such as gain and phase inconsistencies, sensor position biases, and mutual coupling.
-These imperfections introduce additional “array imperfection noise”, which affects the received signal and captures the complexities of practical scenarios.
+Figure 2: The framework of TransDOA and transfer learning approach proposed for array imperfections.
+In this work, two models of array receiving signals are considered. The first is an ideal array signal model, while the second incorporates array imperfections to reflect real-world conditions.
 II-A Ideal Signal Model
@@
 $\boldsymbol{s}(t)={{[{s_{1}}(t),\ldots,{s_{K}}(t)]}^{T}}$ and $\boldsymbol{n}(t)$ denote the transmitted signal and the noise vector at time index $t$, respectively.
-$\boldsymbol{s}(t)$ is Gaussian distributed in both its real and imaginary parts, and $\boldsymbol{n}(t)$ is additive Gaussian noise, with no correlation between any of the signals or noise.
-For a ULA, $\boldsymbol{a}(\boldsymbol{p},\theta_{k})$ can be expressed as:
-$\boldsymbol{a}\left(\boldsymbol{p},\theta_{k}\right)=\left[1,e^{-j\frac{2\pi d}{\lambda}\sin\left(\theta_{k}\right)},\ldots,e^{-j\frac{2\pi d}{\lambda}(M-1)\sin\left(\theta_{k}\right)}\right]^{T}$ (2)
-where $\lambda=c/f$ denotes the wavelength of the incident signals, $d$ is the distance between the array elements, while $c$ and $f$ represent the propagation speed and the frequency of the incident signals, respectively.
-Similarly, for a UCA, $\boldsymbol{a}(\boldsymbol{p},\phi_{k},\theta_{k})$ can be represented in the following form:
-$\boldsymbol{a}\left(\boldsymbol{p},\phi_{k},\theta_{k}\right)=$ $\left[e^{-j\left(2\pi/\lambda\right)\cos(\phi_{1}-\theta_{k})\sin(\phi_{k})},\dots,\right.$ (3)
-$\left.e^{-j\left(2\pi/\lambda\right)\cos(\phi_{M}-\theta_{k})\sin(\phi_{k})}\right]^{T}$
-where $\phi_{m}=\frac{2\pi m}{M}$ denotes the azimuth of the $m$th sensor of the UCA. Based on the received signal vector discussed above, the spatial covariance matrix (SCM) is computed as follows:
-[…]
-where $\mathbb{E}\{\cdot\}$ denotes the expectation operator, ${{\boldsymbol{R}}_{s}}=\mathbb{E}\left\{\boldsymbol{s}(t){{\boldsymbol{s}}^{\mathrm{H}}}(t)\right\}$ represents the covariance matrix of incident signals and $\sigma_{N}^{2}$ stands for the noise power.
-Depending on the statistical degrees of freedom of SCM, $K\leq M$ different DOAs can be estimated precisely.
-However, since the number of snapshots is limited in practice, the SCM is replaced by the sampled SCM calculated via:
+The steering vectors for ULA and UCA can be expressed in the following form:
+$\boldsymbol{a}(\boldsymbol{p},\theta_{k})=[1,e^{-j\frac{2\pi d}{\lambda}\sin(\theta_{k})},\ldots,e^{-j\frac{2\pi d}{\lambda}(M-1)\sin(\theta_{k})}]^{T}$ (2)
+$\boldsymbol{a}\left(\boldsymbol{p},\theta_{k},\phi_{k}\right)=$ $\left[e^{-j\frac{2\pi R}{\lambda}\cos(\phi_{1}-\theta_{k})\sin(\phi_{k})},\dots,\right$. (3)
+$\left.e^{-j\frac{2\pi R}{\lambda}\cos(\phi_{M}-\theta_{k})\sin(\phi_{k})}\right]^{T}$
+where $\theta_{k}$ and $\phi_{k}$ denote the DOA of the $k$-th signal.
+$\lambda=c/f$ denotes the wavelength of the incident signals, $d$ is the distance between the array elements in ULA and $R$ is the radius of the UCA.
+$\phi_{m}=\frac{2\pi m}{M}$ denotes the azimuth of the $m$th sensor of the UCA.
+Based on the received signal discussed above, the sampled covariance matrix (SCM) is computed as follows:
 $\widetilde{\boldsymbol{R}}_{y}=\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{y}(t)\boldsymbol{y}^{\mathrm{H}}(t)$ (4)
-$\widetilde{\boldsymbol{R}}_{y}$ is an unbiased estimate of ${\boldsymbol{R}}_{y}$ calculated by the received signal vector.
-Most existing DOA estimation algorithms focus on deriving the source DOAs $\boldsymbol{\theta}=[{{\theta}_{1}},\ldots,{{\theta}_{K}}]$ from the sampled SCM $\widetilde{\boldsymbol{R}}_{y}$, which has been proven to be the most effective approach.
+$\widetilde{\boldsymbol{R}}_{y}$ is an unbiased estimate of ideal covariance matrix calculated by the received signal vector.
+Most existing DOA estimation algorithms focus on deriving the source DOAs from the sampled SCM $\widetilde{\boldsymbol{R}}_{y}$, which has been proven to be an effective approach.
 II-B Signal Model with Array Imperfection
-In practical applications, the array signal reception model is often not as ideal as its theoretical counterpart.
-Position biases may be introduced during the arrangement of array elements.
-Amplitude and phase inconsistencies may be caused by defects in the design and manufacturing of sensors.
-And mutual coupling effects may be induced during the signal reception process.
+In practical applications, the array signal reception model is often deviates from its theoretical counterpart.
+Such non-ideal conditions typically arise from three primary sources: position biases introduced during array element arrangement; amplitude and phase inconsistencies caused by design and manufacturing defects; and mutual coupling effects induced during the signal reception process.
 In [35], Liu et al.
-established a simplified model to show how these three array imperfections affect the signal model in ULA, and introduced a hyperparameter $\rho\in[0,1]$ to succinctly control the intensity of these array imperfections.
-$\boldsymbol{e}_{\text{pos}}&=\rho\times\left[0,-0.2,\dots,-0.2,0.2,0.2,\dots,0.2\right]^{T}\times d\\ \boldsymbol{e}_{\text{gain}}&=\rho\times\left[0,0.2,\dots,0.2,-0.2,\dots,-0.2\right]^{T}\\ \boldsymbol{e}_{\text{phase}}&=\rho\times\left[0,-30^{\circ},\dots,-30^{\circ},30^{\circ},\dots,30^{\circ}\right]^{T}\\ \boldsymbol{e}_{\text{mc}}&=\rho\times\left[\gamma^{0},\gamma^{1},\dots,\gamma^{M-1}\right]^{T}\end{aligned}$
+established a simplified model to demonstrate how these three array imperfections affect the signal model in a ULA, introducing a hyperparameter $\rho\in[0,1]$ to succinctly control the intensity of these effects.
+$\begin{aligned} \boldsymbol{e}_{\text{pos}}&=\rho\times\left[0,-0.2,\dots,-0.2,0.2,0.2,\dots,0.2\right]^{T}\times d\\ \boldsymbol{e}_{\text{gain}}&=\rho\times\left[0,0.2,\dots,0.2,-0.2,\dots,-0.2\right]^{T}\\ \boldsymbol{e}_{\text{phase}}&=\rho\times\left[0,-30^{\circ},\dots,-30^{\circ},30^{\circ},\dots,30^{\circ}\right]^{T}\\ \boldsymbol{e}_{\text{mc}}&=\rho\times\left[\gamma^{0},\gamma^{1},\dots,\gamma^{M-1}\right]^{T}\end{aligned}$
 where $\gamma=0.3e^{j60^{\circ}}$ is the mutual coupling coefficient between adjacent sensors.
-The term $\boldsymbol{e}_{\text{pos}}$ represents the positional bias of each array element, while $\boldsymbol{e}_{\text{gain}}$ and $\boldsymbol{e}_{\text{phase}}$ denote the amplitude and phase biases, respectively.
-$\boldsymbol{e}_{\text{mc}}$ models the magnitude of the mutual coupling effect for a ULA, where the strength of the mutual coupling is proportional to the square of the distance.
-The matrix $\boldsymbol{E}_{\text{mc}}$, denoted as a Toeplitz matrix with parameter vector $\boldsymbol{e}_{\text{mc}}$ [59], can be employed to model the mutual coupling across the entire array:
+The vectors $\boldsymbol{e}_{\text{pos}}$, $\boldsymbol{e}_{\text{gain}}$ and $\boldsymbol{e}_{\text{phase}}$ denote the element positional, amplitude, and phase biases, respectively.
+Specifically, $\boldsymbol{e}_{\text{mc}}$ models the magnitude of the mutual coupling effect in a ULA, where the strength of the mutual coupling is proportional to the square of the distance between elements.
+The overall mutual coupling across the entire array is represented by the matrix $\boldsymbol{E}_{\text{mc}}$, which is defined as a Toeplitz matrix parameterized by the vector $\boldsymbol{e}_{\text{mc}}$ [44]:
 $\small\boldsymbol{E}_{\text{mc}}=\begin{bmatrix}\gamma^{0}&\gamma^{1}&\cdots&\gamma^{i}&\gamma^{i+1}&\cdots&\gamma^{M-1}\\ \gamma^{1}&\gamma^{0}&\gamma^{1}&\cdots&\gamma^{i}&\cdots&\vdots\\ \vdots&\gamma^{1}&\gamma^{0}&\ddots&\cdots&\ddots&\vdots\\ \gamma^{i}&\cdots&\ddots&\ddots&\ddots&\ddots&\gamma^{i}\\ \gamma^{i+1}&\ddots&\cdots&\ddots&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&\cdots&\ddots&\gamma^{0}&\gamma^{1}\\ \gamma^{M-1}&\cdots&\gamma^{i+1}&\gamma^{i}&\cdots&\gamma^{1}&\gamma^{0}\end{bmatrix}$ (6)
@@
 The steering vector affected by all three array imperfections can be expressed as:
-$\boldsymbol{a}(\theta,\boldsymbol{e})&=\left(\boldsymbol{I}_{M}+\delta_{mc}\boldsymbol{E}_{mc}\right)\times\left(\boldsymbol{I}_{M}+\operatorname{Diag}\left(\delta_{gain}\boldsymbol{e}_{gain}\right)\right)\\ &\times\operatorname{Diag}\left(\exp\left(j\delta_{\text{phase}}\boldsymbol{e}_{\text{phase}}\right)\right)\times\boldsymbol{a}\left(\theta,\boldsymbol{p}+\delta_{pos}\boldsymbol{e}_{pos}\right)\end{aligned}$
-where $\delta(\cdot)$ is used to denote specific array imperfections, $I_{M}$ represents the $M\times M$ identity matrix, while $\operatorname{Diag}(\cdot)$ denotes a diagonal matrix formed from a given vector.
-And $\boldsymbol{a}\left(\theta,\boldsymbol{p}+\boldsymbol{e}_{pos}\right)$ refers to the array steering vector obtained based on the position biases:
+$\begin{aligned} \boldsymbol{a}(\theta,\boldsymbol{e})&=\left(\boldsymbol{I}_{M}+\delta_{mc}\boldsymbol{E}_{mc}\right)\times\left(\boldsymbol{I}_{M}+\operatorname{Diag}\left(\delta_{gain}\boldsymbol{e}_{gain}\right)\right)\\ &\times\operatorname{Diag}\left(\exp\left(j\delta_{\text{phase}}\boldsymbol{e}_{\text{phase}}\right)\right)\times\boldsymbol{a}\left(\theta,\boldsymbol{p}+\delta_{pos}\boldsymbol{e}_{pos}\right)\end{aligned}$
+where $\delta(\cdot)$ indicates the presence of specific array imperfections.
+$\boldsymbol{I}_{M}$ represents the $M\times M$ identity matrix, while $\operatorname{Diag}(\cdot)$ denotes a diagonal matrix formed from a given vector.
+And $\boldsymbol{a}\left(\theta,\boldsymbol{p}+\boldsymbol{e}_{pos}\right)$ refers to the array steering vector incorporating the position biases:
 $\boldsymbol{a}\left(\theta_{k},\boldsymbol{p}+\boldsymbol{e}_{pos}\right)=$ $[e^{-j\frac{2\pi}{\lambda}(0d+e_{\text{pos}}^{0})\sin(\theta_{k})},e^{-j\frac{2\pi}{\lambda}(1d+e_{\text{pos}}^{1})\sin(\theta_{k})},$ (8)
@@
 $\boldsymbol{A}(\boldsymbol{\theta},\boldsymbol{e})=\left[\boldsymbol{a}(\theta_{1},\boldsymbol{e}),\boldsymbol{a}(\theta_{2},\boldsymbol{e}),\ldots,\boldsymbol{a}(\theta_{K},\boldsymbol{e})\right]$ (9)
-By substituting $\boldsymbol{A}(\boldsymbol{\theta},\boldsymbol{e})$ for the steering matrix $\boldsymbol{A}(\boldsymbol{p},\boldsymbol{\theta})$ in Section II-A, a more accurate simulation of the array signal reception model in practical scenarios can be achieved.
+By substituting $\boldsymbol{A}(\boldsymbol{\theta},\boldsymbol{e})$ for the steering matrix $\boldsymbol{A}(\boldsymbol{p},\boldsymbol{\theta})$ introduced in Section II-A, a more accurate simulation of the array signal reception model in practical scenarios can be achieved.
 III NETWORKS FOR DOA ESTIMATION
-Figure 2: The overrall structure of DOA-ViT and the processing procedure of input covariance matrix.
-Given the remarkable performance of the Vision Transformer (ViT) model across various domains, its strong nonlinear representation capability and efficient feature learning ability are expected to enhance subsequent transfer learning.
-Therefore, we adopt ViT as our baseline model.
-The overall architecture of the ViT model designed for DOA estimation, referred to as DOA-ViT, is illustrated in Figure 2.
-In terms of model design, we adhere to the original Vision Transformer framework [60] as closely as possible to ensure consistency and leverage its proven architecture.
+Figure 3: The overall structure of TansDOA and the processing procedure of input covariance matrix.
+To capture global dependencies among array elements, we propose a Transformer-based model named TransDOA.
+The overall framework comprising the TransDOA architecture and the transfer learning strategy is illustrated in Fig.
+2.
+Specifically, the detailed structure of the TransDOA is further presented in Fig.
+3.
 III-A Covariance Matrix Embedding
-Since the transformer encoder can only operate on a set of input vectors, it is necessary to encode the input covariance matrix into a set of vectors.
-To achieve this, the split and vectorization operation is employed, which is detailed in (11) and (12).
-The sampled SCM $\widetilde{\boldsymbol{R}}_{y}$ is partitioned column-wise into $M$ smaller blocks and subsequently each block is flattened into a vector of size $2M$, resulting in $M$ vectors.
-Finally, a linear layer is utilized to encode each input vector, generating a matrix of dimensions $M\times D$, which represents $M$ vectors, each of dimension $D$.
+To accommodate the Transformer’s sequence-based architecture for DOA estimation, the SCM $\widetilde{\boldsymbol{R}}_{y}$ is partitioned and mapped into a sequence of vectors.
+Specifically, $\widetilde{\boldsymbol{R}}_{y}$ is split column-wise into $M$ blocks, each flattened and transformed into a $D$-dimensional embedding $\boldsymbol{x}_{i}$ via a linear layer $\boldsymbol{E}$, as formulated in (10)–(12).
+A learnable DOA token is prepended to the input sequence to aggregate global features for the final estimation task (13).
 $\widetilde{\boldsymbol{R}}_{y}$ $=\left[\boldsymbol{r}_{1},\boldsymbol{r}_{2},\boldsymbol{r}_{3},\ldots,\boldsymbol{r}_{M}\right],\;\small{\widetilde{\boldsymbol{R}}_{y}\in\mathbb{C}^{M\times M},r_{i}\in\mathbb{C}^{M}}$ (10)
-$\widetilde{\boldsymbol{r}}_{i}$ $=\left[\text{real}(\boldsymbol{r}_{i});\text{imag}(\boldsymbol{r}_{i})\right],\quad\widetilde{\boldsymbol{r}}\in\mathbb{R}^{2M}$ (12)
+$\widetilde{\boldsymbol{r}}_{i}$ $=\left[\text{real}(\boldsymbol{r}_{i});\text{imag}(\boldsymbol{r}_{i})\right],\quad\widetilde{\boldsymbol{r}}_{i}\in\mathbb{R}^{2M}$ (11)
 $\boldsymbol{x}_{i}$ $=\boldsymbol{E}\widetilde{\boldsymbol{r}}_{i},\quad\boldsymbol{E}\in\mathbb{R}^{D\times 2M}$ (12)
 $\boldsymbol{X}_{\text{input }}$ $=\left[doa\;token,\boldsymbol{x}_{1},\boldsymbol{x}_{2},\ldots,\boldsymbol{x}_{M}\right]$ (13)
-The formulation of Covariance Matrix Embedding procedure is presented in Eqs.
-(11) to (14).
-Notably, similar to the widely used transformer models, a learnable $D$-dimensional vector, called the DOA estimation token, is added to the first column of the embeded matrix.
-Since the transformer encoder captures global dependencies through self-attention, the DOA estimation token effectively integrates information from the entire sequence and is subsequently used in the final task, such as classification and regression.
-This approach simplifies model design and enables the transformer model to directly leverage global context for task prediction.
-III-B Position Embedding
-Transformer models inherently lack information about the order of input vectors.
-In other words, even if the positions of input vectors are randomly shuffled, the transformer encoder produces identical outputs for each input vector, as it treats the input vectors as an unordered set.
-To address this limitation, positional embeddings are introduced to encode the relative or absolute positions of the input vectors.
-Specifically, a set of learnable vectors, known as positional embeddings, is added directly to the input vectors as additional positional encoding information:
+Since the self-attention mechanism is permutation-invariant, learnable positional embeddings $\boldsymbol{E}_{pos}$ are added to the input representations to preserve the spatial structure of the array signal:
 $\boldsymbol{z}_{\text{input }}$ $=\left[\boldsymbol{x}_{0}+\boldsymbol{e}_{0},\boldsymbol{x}_{1}+\boldsymbol{e}_{1},\cdots,\boldsymbol{x}_{M}+\boldsymbol{e}_{M}\right]$ (14)
 $=\boldsymbol{X}_{\text{input }}+\boldsymbol{E}_{\text{pos }},\quad\boldsymbol{E}_{\text{pos }}\in\mathbb{R}^{D\times(M+1)}$
-By adding position embeddings to the input representations, the model can incorporate positional information into its computations, ensuring that the order of the input vectors influences the final representation.
-This mechanism is essential for tasks where understanding the sequential arrangement of input vectors significantly impacts the model’s performance, such as DOA estimation.
-III-C Transformer Encoder
-Figure 3: The framework of transfer learning approach proposed for array imperfections.
-The Multi-Head Self-Attention (MHSA) mechanism is a key component of the Transformer Encoder, which is designed to enhance the model’s capacity to capture complex relationships within sequences.
-Given an input $\boldsymbol{Z}\in\mathbb{R}^{D\times(M+1)}$, where $D$ is the feature dimension and $M+1$ is the sequence length, the MHSA operates are summarized as follow.
+Further details regarding the SCM Embedding process and the comprehensive configuration of the TransDOA model are presented in Fig. 3.
+III-B Transformer Encoder
+The Multi-Head Self-Attention (MHSA) mechanism serves as the core component of the Transformer Encoder, designed to capture complex global dependencies within the input sequence.
+Given an input $\boldsymbol{Z}\in\mathbb{R}^{D\times(M+1)}$, where $D$ is the feature dimension and $M+1$ is the sequence length, the MHSA operates are defined as follow.
 Linear Projections: The input is linearly projected into queries $\boldsymbol{Q}$, keys $\boldsymbol{K}$, and values $\boldsymbol{V}$ for each attention head $i$:
@@
 As illustrated in Fig.
-2, the Transformer Encoder, consisting of layer normalization, MHSA and an MLP, enables the DOA-ViT model to effectively capture and integrate global information.
-III-D DOA Estimation Head
-The DOA Estimation Head serves as the final layer for DOA estimation.
-In this module, the vector corresponding to the DOA estimation token is extracted; then a linear layer is directly applied to regress the target DOA values.
-The loss function between the predicted DOA values and the true DOA is expressed as:
+3, the Transformer Encoder, consisting of layer normalization, MHSA and an MLP, enables the TransDOA model to effectively capture and integrate global information.
+III-C DOA Estimation Head
+The DOA Estimation Head serves as the final regression module.
+The feature vector corresponding to the aggregated DOA token is extracted and projected via a linear layer to predict the target DOAs.
+To address the permutation ambiguity inherent in multi-source estimation, we employ the permutation invariant training loss:
 $l(\boldsymbol{\widetilde{\boldsymbol{R}}},\boldsymbol{\theta};\boldsymbol{\psi})=\min_{\boldsymbol{P}\in\mathcal{P}}\left(\frac{1}{K}\left\|\boldsymbol{\theta}-\boldsymbol{P}\hat{\boldsymbol{\theta}}(\widetilde{\boldsymbol{R}};\boldsymbol{\psi})\right\|^{2}\right)^{\frac{1}{2}}$ (18)
-where $\boldsymbol{\psi}$ represents all the parameters of the model and $\hat{\boldsymbol{\theta}}$ is the estimation result of the DOA-ViT model, which includes $K$ DOAs.
-$\boldsymbol{\theta}$ corresponds to the true DOAs, and the set $\mathcal{P}$ encompasses all $M\times M$ permutation matrices, defined as binary matrices where each row and column contains exactly one non-zero element.
-The loss function for a dataset $\mathcal{D}$ is described in the following manner:
-$\mathcal{L}_{\mathcal{D}}(\boldsymbol{\psi})=\frac{1}{|\mathcal{D}|}\sum_{(\boldsymbol{\theta}^{(j)},\widetilde{\boldsymbol{R}}^{(j)})\in\mathcal{D}}l(\widetilde{\boldsymbol{R}}^{(j)},\boldsymbol{\theta}^{(j)};\boldsymbol{\psi})$ (20)
-where $|\mathcal{D}|$ denotes the number of samples in the simulated dataset and $\widetilde{\boldsymbol{R}}^{(j)}$ is the $j$th sampled SCM in the dataset.
-It is worth noting that our proposed method can be easily extended to arbitrary array configurations and 2D (Two-Dimensional) DOA estimation.
-If one aims to implement 2D DOA estimation, the DOA Estimation Head can be directly extended to predict the values of $\theta$ and $\phi$.
-The corresponding loss function for this approach is derived in the following form:
+where $\boldsymbol{\psi}$ denotes the model parameters and $\hat{\boldsymbol{\theta}}$ represents the estimation result, which includes $K$ DOAs.
+$\boldsymbol{\theta}$ corresponds to the true DOAs, and the set $\mathcal{P}$ encompasses all $K\times K$ permutation matrices, defined as binary matrices where each row and column contains exactly one non-zero element.
+Furthermore, the proposed framework extends naturally to 2D (Two-Dimensional) DOA estimation.
+By extending the DOA Estimation Head to predict both azimuth $\theta$ and elevation $\phi$ angles, the joint loss function is formulated as:
 $\begin{split}l(\boldsymbol{\widetilde{\boldsymbol{R}}},\boldsymbol{\theta},\boldsymbol{\phi};\boldsymbol{\psi})=&\min_{\boldsymbol{P}\in\mathcal{P}}\left(\frac{1}{2K}\left\|\boldsymbol{\theta}-\boldsymbol{P}\hat{\boldsymbol{\theta}}(\widetilde{\boldsymbol{R}};\boldsymbol{\psi})\right\|^{2}\right.\\ &\quad\left.+\frac{1}{2K}\left\|\boldsymbol{\phi}-\boldsymbol{P}\hat{\boldsymbol{\phi}}(\widetilde{\boldsymbol{R}};\boldsymbol{\psi})\right\|^{2}\right)^{\frac{1}{2}}\end{split}$ (19)
 IV Transfer Learning Approach
-By leveraging large-scale datasets for training, deep neural networks have achieved superior performance in DOA estimation.
-However, due to the statistical differences in the sampled SCM between simulated data and real-world data with array imperfections, the model’s performance in practical scenarios is adversely affected.
-Inspired by recent advancements in transfer learning [61, 62], and in order to address the significant performance degradation of deep learning models due to array imperfections in practical scenarios, we propose a supervised transfer learning approach, which is designed to enhance the adaptability of deep learning methods to array imperfections.
-IV-A Goal of Transfer Learning
-Transfer learning typically involves two tasks: the source task and the target task.
-The source task usually has abundant data available to effectively train the model, whereas the target task has a limited amount of data.
-Building on machine learning methods, transfer learning allows prior knowledge from the source task to be applied to the target task, even though there are different data distributions between the source and target tasks [55].
-IV-B Source and Target Domains Modeling
-In our transfer learning approach, two distinct scenarios are considered.
-The first is the ideal array signal reception scenario, which serves as a scenario that most deep learning models focus on.
-In this scenario, training data can typically be generated in large quantities through simulation programs, resulting in ample data availability at a relatively low cost.
-Consequently, deep learning algorithms can be effectively trained and validated based on the extensive dataset, ultimately achieving satisfactory performance.
-The second scenario pertains to practical applications, where array imperfections in the antenna system are present.
-In this case, data should be obtained through actual measurements, leading to higher costs and a relatively limited dataset.
-The ideal scenario data is treated as source domain data, and the practical scenario data as target domain data.
-Therefore, how to effectively leverage the prior knowledge acquired from a model trained on a substantial amount of simulation data to enhance model’s performance in a data-scarce or even data-absent practical setting becomes a valuable area for exploration.
-IV-C Transfer learning Framework
-Assume that we have already trained a model with weight parameters $\boldsymbol{\psi}_{s}$ that performs well on simulated data.
-This model is capable of effectively extracting features and performing DOA estimation in the source domain.
-However, due to the presence of array imperfections, substantial statistical discrepancies arise between the data distributions of the source and target domains.
-Consequently, when applied to the target domain, the model with parameters $\boldsymbol{\psi}_{s}$ extracts features that are significantly influenced by these statistical discrepancies, thereby deteriorating DOA estimation performance.
-To mitigate this issue, the objective of this work is to train a feature extractor for the target domain that is designed to learn domain-invariant features, which remain unaffected by domain shifts induced by array imperfections, such as element position biases, amplitude-phase inconsistencies and mutual coupling effects that distort the statistical properties of the sampled SCM.
-Only the essential information relevant to the DOA estimation task is expected to be retained in domain-invariant features, thereby enhancing the model’s generalization ability in practical scenarios.
-The framework of our transfer learning framework is illustrated in Fig.
-II, and we will provide a detailed introduction in the following.
-In our transfer learning approach, we remove the DOA estimation Head from the DOA-ViT model and treat the remaining structure as a feature extractor, denoted as $\mathcal{F}$, which allows us to extract features from both the source and target domains:
+Figure 4: The framework of transfer learning approach proposed for array imperfections.
+Deep learning models trained on large-scale simulated datasets often suffer from performance degradation in practical scenarios. This is primarily due to the domain gap caused by array imperfections.
+To bridge this gap, we propose a supervised transfer learning framework based on feature alignment.
+As illustrated in Fig.
+III, our approach aims to train a target domain feature extractor that maps distorted SCM into a domain-invariant feature space aligned with the ideal array manifold.
+IV-A Domain Characterization
+In our approach, two distinct scenarios are considered.
+The first is the ideal reception scenario, a standard paradigm in deep learning research.
+In this scenario, large-scale training datasets can be generated via simulation at low cost, enabling effective model training and superior performance.
+The second scenario pertains to practical applications characterized by array imperfections.
+In contrast to the ideal case, data acquisition here relies on physical measurements, inevitably resulting in high costs and data scarcity.
+Based on these characteristics, we formally define the respective domains as follows:
+  • Source Domain ($\mathcal{D}_{s}$): Consists of abundant ideal SCMs generated via simulation. A model trained here captures the theoretical array manifold but fails on imperfect data.
+  • Target Domain ($\mathcal{D}_{t}$): Consists of limited labeled data from the practical system with array imperfections.
+How to effectively leverage the prior knowledge acquired from a model trained on a substantial amount of simulation data to enhance model’s performance in a data-scarce or even data-absent practical setting becomes a valuable area for exploration.
+IV-B Transfer learning Framework
+To address the performance degradation caused by statistical discrepancies between the source and target domains, we propose a transfer learning framework, as illustrated in Fig.
+III.
+By leveraging the pre-trained source parameters $\boldsymbol{\psi}_{s}$, our objective is to train a target extractor that learns domain-invariant features, thereby ensuring robustness against array imperfections.
+Specifically, we remove the DOA estimation Head from the TransDOA model and utilize the backbone as a feature extractor, denoted as $\mathcal{F}$.
+The feature representations for the source and target domains are obtained as:
 $\boldsymbol{z}_{s,i}$ $=\mathcal{F}_{s}(\widetilde{\boldsymbol{R}}_{s}^{(i)};\boldsymbol{\psi}_{s})$ (20)
 $\boldsymbol{z}_{t,i}$ $=\mathcal{F}_{t}(\widetilde{\boldsymbol{R}}_{t}^{(i)};\boldsymbol{\psi}_{t})$
-where $\mathcal{F}_{s}$ and $\mathcal{F}_{t}$ represent the feature extractors for the source and target domains, respectively, while $\boldsymbol{\psi}_{s}$ and $\boldsymbol{\psi}_{t}$ represent the model parameters for the source and target domain models.
-$\boldsymbol{z}_{s,i}$ represents the features extracted from the $i$-th source domain input, and $\boldsymbol{z}_{t,i}$ denotes the features extracted from the $i$-th target domain input.
-In practical scenarios, a dataset can be constructed by collecting a series of sampled SCMs along with their corresponding DOA values.
-Base on the true DOA values, a simulation data generator is employed to generate corresponding ideal covariance matrices that are unaffected by array imperfections, which serve as the source domain data to be aligned.
-As illustrated in Fig.
-II, during the transfer learning approach, data in the source domain is fed into $\mathcal{F}_{s}$ with frozen pre-trained weights to ensure the effectiveness of feature extraction from the source domain, while data in the target domain is fed into a trainable feature extraction module $\mathcal{F}_{t}$.
-In this way, by aligning the features $\boldsymbol{z}_{t,i}$ to $\boldsymbol{z}_{s,i}$, the target domain feature extractor $\mathcal{F}_{t}$ can be guided to extract features that can be directly used for DOA estimation, thereby mitigating the impact of array imperfections.
-Based on the feature alignment procedure described above, two loss functions are proposed for feature alignment as follows.
-MSE Loss: A mean squared error is directly applied between the features of the source and target domains to achieve the alignment of target domain features with source domain features.
+where $\mathcal{F}_{s}$ and $\mathcal{F}_{t}$ are the feature extractors for the source and target domains, parameterized by $\boldsymbol{\psi}_{s}$ and $\boldsymbol{\psi}_{t}$, respectively.
+The vectors $\boldsymbol{z}_{s,i}$ and $\boldsymbol{z}_{t,i}$ represent the feature extracted from the $i$-th pair of source and target inputs.
+To ensure exact feature correspondence between the source and target domains, it is critical to construct a strictly paired dataset.
+A Simulation Data Generator is employed for this purpose.
+In practical scenarios, we collect a series of SCMs along with their corresponding DOA values.
+Subsequently, these true DOA values are fed into the Simulation Data Generator to generate corresponding ideal covariance matrices.
+This process establishes a one-to-one mapping, where each collected real-world sample is paired with a generated ideal counterpart, thereby forming the requisite paired dataset for effective knowledge transfer and model calibration.
+Utilizing the constructed paired dataset, the transfer learning framework is illustrated in Fig.
+III.
+The pre-trained source feature extractor $\mathcal{F}_{s}$ is frozen to serve as a stable reference.
+The target encoder $\mathcal{F}_{t}$ is then trained to align its extracted features $\boldsymbol{z}_{t,i}$ with the reference features $\boldsymbol{z}_{s,i}$.
+This alignment compels $\mathcal{F}_{t}$ to suppress domain-specific distortions caused by array imperfections.
+To enforce such robust alignment, we propose a composite objective function consisting of two terms:
+1) MSE Loss: The Mean Squared Error (MSE) is employed to minimize the Euclidean distance between the feature representations of the two domains:
 $\mathcal{L}_{\text{mse}}(\boldsymbol{z}_{i}^{S\leftrightarrow T})=\|\boldsymbol{z}_{s,i}-\boldsymbol{z}_{t,i}\|^{2}$ (21)
-[…]
-Cosine Similarity Loss: To further align the structure of features between the source and target domains, we treat the features as vectors in a high-dimensional space.
-A cosine similarity loss is proposed to achieve angular alignment between the source domain vectors and the target domain vectors.
-$\mathcal{L}_{\cos}(\boldsymbol{z}_{i}^{S\leftrightarrow T})=\frac{\boldsymbol{z}_{s,i}\cdot\boldsymbol{z}_{t,i}}{\|\boldsymbol{z}_{s,i}\|\cdot\|\boldsymbol{z}_{t,i}\|}$ (24)
-where $S$ and $T$ represent the source and target domains, respectively. Finally, the overall loss function can be described as follows:
+where $S$ and $T$ represent the source and target domains, respectively. The notation $\|\cdot\|$ denotes the $l_{2}$-norm of a vector.
+2) Cosine Similarity Loss: To further enforce structural consistency, we treat features as vectors in a high-dimensional space and align their geometric directions. The cosine loss is formulated as:
+$\mathcal{L}_{\cos}(\boldsymbol{z}_{i}^{S\leftrightarrow T})=1-\frac{\boldsymbol{z}_{s,i}\cdot\boldsymbol{z}_{t,i}}{\|\boldsymbol{z}_{s,i}\|\cdot\|\boldsymbol{z}_{t,i}\|}$ (22)
+By minimizing this loss, we maximize the cosine similarity, ensuring the target features share the same angular direction as the source features.
+Finally, the overall objective function is a weighted sum of the aforementioned losses:
 $\mathcal{L}_{\text{total}}(\boldsymbol{Z}^{S},\boldsymbol{Z}^{T})=\alpha\mathcal{L}_{\text{cos}}(\boldsymbol{Z}^{S},\boldsymbol{Z}^{T})+\beta\mathcal{L}_{\text{mse}}(\boldsymbol{Z}^{S},\boldsymbol{Z}^{T})$ (23)
 where $\boldsymbol{Z}^{S}$ and $\boldsymbol{Z}^{T}$ represent all feature vectors extracted from the source domain and target domain, respectively, while $\alpha$ and $\beta$ are hyper-parameters controlling the effect of the two loss functions.
-Building upon the loss function of the transfer learning approach, the optimization of the target domain model parameters $\boldsymbol{\psi}_{\text{t}}$ is performed using first-order methods, e.g., Adaptive Moment Estimation (Adam) optimizer.
-To provide a more precise depiction of the overall process, the transfer learning procedure of DOA-ViT using Adam is detailed as Algorithm 1.
+To provide a more precise depiction of the overall process, the transfer learning procedure of TransDOA using Adam is detailed as Algorithm 1.
 Algorithm 1 Transfer learning Approach using Adam
-0:  DOA-ViT weights $\boldsymbol{\psi}_{s}$ pretrained on the source domain, hyperparameter $\epsilon$.
+0: TransDOA weights $\boldsymbol{\psi}_{s}$ pretrained on the source domain, hyperparameter $\epsilon$.
 0: Target domain Dataset $\mathcal{D}_{\text{target}}$, learning rate $\mu$, number of batches $B$, epochs number $e_{\max}$.
-1:  Initialize DOA-ViT weights $\boldsymbol{\psi}_{s}$ for source domain model.
-2:  Initialize DOA-ViT weights $\boldsymbol{\psi}_{t}$ for target domain model.
+1: Initialize TransDOA weights $\boldsymbol{\psi}_{s}$ for source domain model.
+2: Initialize TransDOA weights $\boldsymbol{\psi}_{t}$ for target domain model.
 3:for $\text{epoch}=0,1,\dots,e_{\max}$ do
@@
 V SIMULATIONS AND ANALYSES
-As the DOA estimation system is a typical electronic system, the performance of the model in DOA estimation is influenced by a combination of multiple parameters.
-We summarize the primary factors as follows:
-  1. AS1
-    SNR and the number of snapshots.
-These two parameters are the most extensively discussed in DOA estimation fields.
-It is generally accepted that high SNR and a large snapshots lead to an ideal signal environment, thereby enhancing algorithm performance.
-  2. AS2
-    Number of array elements $\boldsymbol{M}$ and number of incident signals $\boldsymbol{K}$. A larger $M$ and a smaller $K$ facilitate more accurate DOA estimation.
-  3. AS3
-    Angular Separation and Range of incident signals.
-Discussions regarding the angular parameters of incident signals are relatively limited; however, these factors significantly impact a model’s performance.
-In general, smaller angular separations result in more similar steering vectors, leading to increasingly difficult DOA estimation.
-This situation poses a considerable challenge to the model’s angular resolution.
-Additionally, the range of incoming signals can affect the DOA estimation performance of specific array types (such as $\theta$ for ULA or $\phi$ for UCA); greater angles of signal incidence lead to reduced angular resolution for these arrays.
-In this section, multiple distinct simulation scenarios were constructed based on the settings of the aforementioned parameters.
-We demonstrate the effectiveness of the proposed model and transfer learning approach through extensive simulation experiments.
+In this section, we evaluate the performance of the proposed scheme through extensive simulations, comprehensively validating its effectiveness and robustness under varying critical factors, specifically SNR, number of snapshots $T$, number of array elements $M$, number of incident signals $K$, and angular separation.
+Comprehensive comparisons are conducted with a wide range of classic and state-of-the-art model-driven methods, including high-resolution subspace methods such as MUSIC and Unity-ESPRIT algorithms, as well as the compressive sensing-based $\ell_{1}$-SVD approach [18, 45], which has strong performance in low SNR and small snapshots scenarios.
+Additionally, a series of deep learning methods and their variants that integrate traditional algorithms are considered, including the spatial spectrum estimation CNN algorithm (SPE-CNN) [34], SubspaceNet [39], ASL-2 [37], and Learning-SPICE [38].
+SPE-CNN is recognized as a well-established deep learning-based method and SubspaceNet can be understood as a subspace algorithm implemented using deep learning networks.
+Furthermore, ASL-2 and Learning-SPICE, which representing the latest algorithms that combine deep learning and compressive sensing, are expected to perform well in limited snapshots scenarios.
 We begin with a detailed description of the experimental setup in Subsection V-A.
-Subsequently, the performance of our proposed model under these setups is reported in Subsection V-B.
-Furthermore, we demonstrate that our proposed model exhibits significant performance degradation in scenarios involving array imperfections, which motivates the introduction of transfer learning techniques to address the challenges posed by array imperfections.
-The simulation results are presented in Subsection V-C.
-A wide range of classic and state-of-the-art algorithms were selected to demonstrate the effectiveness of the proposed model.
-We first selected several classic subspace-based approaches as baseline methods for comparison, such as MUSIC [17] and Unity-ESPRIT algorithms [16].
-Compressive sensing algorithm $\ell_{1}$-SVD algorithm [24, 63] is considerd, which has strong performance in low SNR and small snapshots scenarios.
-In addition, a series of deep learning methods and their variants that integrate traditional algorithms were used for comparative analysis.
-We selected spatial spectrum estimation CNN algorithm (SPE-CNN) [1], SubspaceNet [35], ASL-2 [33], and Learning-SPICE [34] for comparison.
-SPE-CNN is recognized as a well-established deep learning-based method and SubspaceNet can be understood as a subspace algorithm implemented using deep learning networks.
-Furthermore, ASL-2 and Learning-SPICE, which representing the latest algorithms that combine deep learning and compressive sensing, are expected to perform well in small snapshots scenarios.
+The performance of the proposed model under these setups is then presented in Subsection V-B.
+Furthermore, in Subsection V-C, we validate the effectiveness of transfer learning strategies through simulations conducted in practical scenarios involving array imperfections.
 V-A Experimental Setup
-1) Simulation Scenarios: Throughout this study, we consider four distinct scenarios to evaluate the performance of our proposed algorithm:
-  • Scenario 1: A ULA with $M=8$ and $K=3$.
-  • Scenario 2: A ULA with $M=8$ and $K=7$.
-  • Scenario 3: A ULA with $M=16$ and $K=3$.
-  • Scenario 4: A UCA with $M=12$ and $K=5$.
-Figure 4: DOA estimation under Scenarios 1 with $SNR=-5db$ and $snap=10$.
-Three random incident signals are generate and we focus on the first one.
-The points represent the estimated DOA provided by the algorithms while the circles denote the true DOAs.
-TABLE I: CDF-Based Statistical Evaluation results under Scenarios 1 with $snap=10$ and $SNR=-5$.
-Algorithms Ours MUSIC Unity-ESPRIT $\ell_{1}$-SVD SPE-CNN ASL-2 SubspaceNet Learning-SPICE
-$\text{CDF}(3)$ 55.36% 54.10% 48.90% 44.30% 60.11% 33.49% 52.63% 68.08%
-$\text{CDF}(10)$ 89.20% 73.43% 80.31% 82.93% 81.35% 55.68% 87.35% 87.51%
-CDF10% $0.455^{\circ}$ $0.410^{\circ}$ $0.526^{\circ}$ $0.638^{\circ}$ $0.368^{\circ}$ $0.766^{\circ}$ $0.501^{\circ}$ $\mathbf{0.309^{\circ}}$
-CDF90% $\mathbf{10.63^{\circ}}$ $38.13^{\circ}$ $24.22^{\circ}$ $20.91^{\circ}$ $29.66^{\circ}$ $55.20^{\circ}$ $13.55^{\circ}$ $16.94^{\circ}$
-Scenario 1 represents a standard multi-target DOA estimation scenario, which is widely adopted for evaluating model performance.
-Subsequently, to further investigate whether the model’s performance deteriorates when handling a larger number of incident angles, Scenario 2 is introduced.
+TABLE I: Simulation Scenarios and Parameter Configurations
+Scen. Setting $(M,K)$ Configuration & Test Conditions
+1 ULA $(8,3)$ DOA: Uniform, Equidistant
+Params [1]: SNR =5=-5 dB, $T=10$
+Params [2]: SNR $\in[-20,5]$ dB, $T=10$
+Params [3]: $T\in[10,100]$, SNR=10=-10 dB
+2 ULA $(8,7)$ DOA: Uniform
+Params [1]: SNR $\in[-20,5]$ dB, $T=10$
+3 ULA $(16,3)$ DOA: Uniform
+Params [1]: SNR $\in[-20,5]$ dB, $T=10$
+4 UCA $(12,5)$ DOA: Deterministic, Uniform
+Params [1]: SNR =5=-5 dB, $T=50$
+Params [2]: SNR =5=5 dB, $T=10$
+  • Uniform: DOAs are sampled from a uniform distribution, subject to a minimum angular separation of $3^{\circ}$ to avoid source clustering.
+  • Equidistant: DOAs are equally spaced across the entire angular sector.
+  • Deterministic: A predetermined set of specific DOAs.
+1) Simulation Scenarios: To facilitate a holistic assessment of the algorithm’s performance, we conduct simulations across four distinct scenarios, as summarized in Table I.
+Figure 5: DOA estimation results under Scenarios 1 with SNR=5dbSNR=-5db and $snap=10$.
+Three incident signals are generate and we focus on the first one.
+The points represent the estimated DOA while the circles denote the true DOAs.
+TABLE II: CDF-Based Statistical Evaluation results under Scenarios 1 with SNR=5SNR=-5 and $snap=10$.
+Method Miss Prob. $\downarrow$ OSPA Error RMSE () MAE () Acc. (%) ECDF Quantile (Raw) ECDF Quantile (Matched)
+Linear $\downarrow$ Square $\downarrow$ Raw $\downarrow$ Matched $\downarrow$ Raw $\downarrow$ Matched $\downarrow$ Raw $\uparrow$ Matched $\uparrow$ 10% $\downarrow$ 90% $\downarrow$ 10% $\downarrow$ 90% $\downarrow$
+MUSIC 0.001 6.65 9.52 20.39 12.11 10.81 6.68 74.3 80.9 0.43 37.31 0.38 30.00
+Unity-ESPRIT 0.000 6.27 8.46 17.09 10.52 8.92 6.29 78.8 82.8 0.50 26.62 0.43 20.82
+$\ell_{1}$-SVD 0.088 6.35 8.36 14.33 9.36 7.60 5.63 80.8 85.4 0.62 20.04 0.58 12.38
+SPE-CNN 0.000 5.53 8.08 18.79 10.54 8.84 5.55 80.6 84.9 0.39 29.47 0.32 25.99
+Learning-SPICE 0.000 4.47 6.34 15.09 9.07 6.43 4.50 87.4 89.1 0.31 18.98 0.28 13.82
+ASL-2 0.049 11.62 14.28 31.66 15.57 19.91 11.06 54.6 60.7 0.77 60.00 0.62 30.00
+SubspaceNet 0.000 5.15 6.79 12.71 8.88 6.47 5.16 85.8 88.3 0.52 16.30 0.48 12.27
+TransDOA (ours) 0.000 4.68 5.99 9.65 7.73 5.18 4.69 88.2 89.0 0.46 11.60 0.45 10.93
+Scenario 1 establishes a standard benchmark for multi-source estimation for a ULA.
+Scenario 2 considers a high-density source environment where the number of signals is near the degrees of freedom of the array.
 Scenario 3 represents a scenario with more array elements, which aims to achieve precise DOA estimation.
-In such cases, algorithms typically demand significantly greater computational resources.
-Thus, maintaining accuracy while minimizing computational overhead becomes a critical challenge in this scenario.
 Scenario 4 involves the use of a UCA for two-dimensional localization.
-This scenario evaluates the model’s capability to extend from one-dimensional to two-dimensional tasks, and tests its performance in 2D DOA estimation.
-The aforementioned scenarios essentially modify the parameters of AS2.
-Within each scenario, data with varying SNR, snapshot numbers, angular separations and angular ranges are thoroughly tested.
-This comprehensive evaluation aims to provide a thorough assessment of the model’s performance, including model’s adaptability to low SNR and small snapshots scenarios, model’s angular resolution capabilities, and model’s angular estimation range in various scenarios.
-2) Hyperparameter Settings: In the developed DOA-ViT model, the dimensionality of the embedded vectors $D$ is fixed at 768, while the dimension of each head $d_{k}$ is set to 64.
-Additionally, our model consists of 6 Transformer encoders and the MHSA module is configured with 12 attention heads.
-Within each encoder, the expansion factor of the MLP layers is set to 4.
-Under each configuration, 50,000 samples are generated for model training, and 20,000 samples are generated for model validation.
-For a ULA, the incident angles $\theta$ are randomly generated within the range $[-60^{\circ},60^{\circ}]$ in the training and validation sets.
-And for a UCA, the incident angles $\theta$ are randomly generated within $[-180^{\circ},180^{\circ}]$, and $\phi$ values are randomly generated within $[0^{\circ},60^{\circ}]$.
-The Adam optimizer is configured with a learning rate of $0.0001$; betas values $(0.9,0.999)$ are used for the first and second moment estimates, which is considered effective in most scenarios.
-The number of training epochs is set to 500.
-However, if the validation loss does not improve over the course of 30 epochs, early stopping is triggered to halt the training.
-For the sake of simplicity and fairness, we train each deep learning model for comparison using the same dataset and optimizer to ensure consistency in the experiments and comparability of the results.
+This scenario evaluates the model’s capability to extend from one-dimensional to two-dimensional tasks.
+To ensure a comprehensive evaluation, we performed simulations across a range of angular configurations, SNRs, and snapshots in each scenario. The specific parameter settings are detailed in Table I.
+2) Implementation Details: The proposed TransDOA model comprises 6 Transformer encoder layers.
+the embedding dimension $D$ is set to 768 and each MHSA module employs 12 attention heads.
+Within each encoder, the expansion factor of the feed-forward network is set to 4.
+Under each simulation scenario, a total of 50,000 training samples and 20,000 validation samples are generated.
+For the ULA, the DOAs $\theta$ are randomly generated within the range $60^{\circ},60^{\circ}]$ in the training and validation sets.
+And for the UCA, the azimuth angle $\theta$ and elevation angle $\phi$ are randomly generated within $180^{\circ},180^{\circ}]$ and $0^{\circ},60^{\circ}]$, respectively.
+Model optimization is performed using the Adam optimizer with a learning rate of $10^{-4}$ and exponential decay rates $(\beta_{1},\beta_{2})=(0.9,0.999)$.
+The number of training epochs is set to 500, with an early stopping mechanism triggered if the validation loss fails to improve for 30 epochs.
+To ensure a fair and rigorous comparison, all deep learning models are trained using the identical dataset and optimization protocol, thereby maintaining experimental consistency and the comparability of results.
 For more detailed implementation and configuration specifics, the complete code is available in our open-source repository [^1] https://github.com/zzb-nice/DOA_est_Master, providing further insight into the structure and parameter settings employed.
-V-B DOA estimation performance based on ViT model
+V-B DOA estimation performance based on TransDOA model
 V-B1 Simulation results in Scenario 1
-A series of simulation experiments were conducted to validate the effectiveness of the proposed ViT model. First, we evaluate the performance of the model in Scenario 1.
-Visual Performance Demonstration: In order to provide a visual comparison of the performance of various algorithms in Scenario 1, specifically in the low SNR and small snapshot conditions, we visually demonstrate the DOA estimation results of multiple algorithms in Fig 4.
-Scenario 1 with SNR set to -5 dB and snapshots number set to 10 is considered, which serves as the low SNR, small snapshot scenario.
-The absolute angular error of the model’s DOA estimates is plotted on the right side.
-Through a direct comparison, the proposed method demonstrates superior performance, with the majority of estimation errors being smaller than $25^{\circ}$ in such a low SNR, small snapshot scenario.
-In contrast, all other algorithms exhibit severe deviations with absolute errors exceeding 50 degrees.
-Figure 5: RMSE $vs.$ SNR results under Scenarios 1 with $snap=10$. The incident signal’s DOA values are randomly generated with the minimum angular interval set to 3 degrees.
-RMSE $\boldsymbol{vs.}$ SNR: To effectively evaluate the model’s performance, it is essential to assess its statistical performance.
-We demonstrate the Root Mean Squared Error (RMSE) of estimated DOA values under varying SNR and snapshots (AS1).
-In Fig.
-5, we investigate the RMSE vs.
-SNR results based on 5,000 simulations.
-It can be seen that the proposed method achieves the best performance when the incident angles are randomly generated, which can be regarded as the most general case.
+To validate the effectiveness of the proposed TransDOA model, we first evaluate its performance in Scenario 1 through a series of targeted experiments.
+Visual Performance Demonstration: To visually evaluate the robustness of various algorithms under challenging signal environments, Fig.
+5 illustrates the DOA estimation results in Scenario 1 with SNR= -5 dB and TT=10, which serves as the low SNR, limited snapshot scenario.
+The right-hand plot in the figure depicts the absolute angular error of each method.
+As illustrated, the proposed method demonstrates superior performance, with most estimation errors remaining below $25^{\circ}$, whereas all other algorithms suffer from severe deviations, with absolute errors exceeding 50 degrees.
+Figure 6: RMSE $vs.$ SNR results under Scenarios 1 with Uniform configuration and Params [2] specified in Table I.
+RMSE $\boldsymbol{vs.}$ SNR: To evaluate the statistical performance of the model, we report the Root Mean Squared Error (RMSE) of the estimated DOAs under varying SNR conditions based on 5,000 simulations.
+As illustrated in Fig.
+6, The proposed method achieves the best RMSE performance when DOAs are randomly generated (Uniform configuration), which can be regarded as the most general case.
 Learning-SPICE closely approaches the performance of the proposed model as SNR increase ($SNR=0db$ or $SNR=5db$).
@@
 $\ell_{1}$-SVD exhibits consistently stable performance, and MUSIC and ASL-2 exhibit inferior performance compared with the aforementioned algorithms.
-Furthermore, in order to provide a more intuitive demonstration of the model’s statistical performance and a more comprehensive evaluation of its overall performance, the cumulative distribution function (CDF) of the DOA estimation absolute error is illustrated to assess the model’s performance, which describes the probability distribution of the absolute error.
+To provide a more granular evaluation of the statistical characteristics, the empirical cumulative distribution function (ECDF) of the absolute DOA estimation error is illustrated to assess the model’s performance.
 As shown in Fig.
-6, 5,000 simulations are conducted in each condition, where the incident signal’s DOA values are randomly generated with the minimum angular interval set to 3 degrees.
-Through the CDF, it can be observed that under higher SNR conditions (e.g., $SNR=0db$ or $5db$), the proposed model, Unity-ESPRIT, SPE-CNN and Learning-SPICE exhibited commendable performance.
-However, compared with other algorithms, the proposed model exhibits a lower cumulative probability of small absolute errors, and it shows a distinct advantage in larger absolute error ranges.
-This trend is more evident in low SNR scenarios like $SNR\leq 0db$, as detailed in the following analysis.
+7, The ECDF provides insight into the error distribution across 5,000 simulations under Scenarios 1 with Uniform configuration.
+As observed from the ECDF curves, while the proposed model exhibits a slightly lower cumulative probability within the high-precision regime, it demonstrates a distinct advantage in larger error ranges.
+This indicates that the proposed method prioritizes global robustness, making it significantly less prone to catastrophic estimation failures.
+This trend is particularly evident in low SNR scenarios (e.g., $SNR\leq 0db$).
 (a) $SNR=0db$
 (b) $SNR=5db$
-(c) $SNR=-5db$
-(d) $SNR=-10db$
-Figure 6: CDF of the absolute error under Scenarios 1 with $snap=10$ and various $SNR$, for (a) $SNR=0db$, (b) $SNR=5db$, (c) $SNR=-5db$ and (d) $SNR=-10db$.
+(c) SNR=5dbSNR=-5db
+(d) SNR=10dbSNR=-10db
+Figure 7: ECDF of the absolute error under Scenarios 1 with with Uniform configuration and Params [2], for (a) $SNR=0db$, (b) $SNR=5db$, (c) SNR=5dbSNR=-5db and (d) SNR=10dbSNR=-10db.
 A representative scenario is depicted in Fig.
-6, where it can be observed that the distribution of absolute errors within the range of $0^{\circ}$ to $3^{\circ}$ for the proposed model is slightly lower than those of Learning-SPICE, SPE-CNN and MUSIC, while exhibiting advantage over the SubspaceNet, Unity-ESPRIT, $\ell_{1}$-SVD and ASL-2 algorithms.
-However, the proposed model’s probability of absolute errors exceeding $10^{\circ}$ is substantially lower, which indicates a reduced frequency of high-error occurrences, highlighting our model’s superior robustness relative to other approaches.
-Figure 7: RMSE $vs.$ SNR results under Scenarios 1 with $snap=10$.
-The dashed line in the figure represents Scene 1, while the solid line represents Scene 2.
-In Scene 1, the true DOA values are [10°, 13°, 16°], and in Scene 2, the true DOA values are [10°, 20°, 30°].
-Figure 8: RMSE $vs.$ SNR results under Scenarios 2 with $snap=10$. The incident signals are randomly generated with the minimum angular interval set to 3 degrees.
-Figure 9: RMSE $vs.$ snap results under Scenarios 1 with $SNR=-10db$.
-The dashed line in the figure represents Scene 1, while the solid line represents Scene 2.
-In Scene 1, the true DOA values are [5.1°, 30.3°, 54.5°], and in Scene 2, the true DOA values are [-9.9°, 15.3°, 39.5°].
-Figure 10: RMSE $vs.$ SNR results under Scenarios 3 with $snap=10$. The incident signals are randomly generated with the minimum angular interval set to 3 degrees.
-Furthermore, through the CDF, a series of statistical measures related to the absolute error can be introduced.
-For example, CDF50% represents the absolute error corresponding to a cumulative probability of ${50\%}$.
-As a quantile-based measure of absolute error, it differs from the commonly used mean absolute error (MSE), which represents the mean of absolute errors.
-Both serve as statistical measures of absolute error and hold statistical significance.
-Similarly, CDF(5) refers to the probability that the absolute error is less than 5, which can also reflect the model’s performance.
-The results of the CDF-based statistical evaluation for various models are presented in Table I.
-From the table, it can be observed that Learning-SPICE demonstrates superior performance at $\text{CDF}(3)$ and CDF10%, achieving the highest DOA estimation accuracy, where the algorithm’s estimated values are most likely to be close to the true values.
-However, our proposed model demonstrates the best performance at $\text{CDF}(10)$ and CDF90%, showing greater robustness by reducing the occurrence of large errors.
-Due to the fact that RMSE penalizes larger estimation errors severely, the proposed model demonstrates superior performance in terms of RMSE compared with the Learning-SPICE algorithm.
-To further demonstrate the performance of the proposed model, we selected four representative models mentioned above for testing in smaller angular interval scenarios.
-We consider two Scenes with different angular separations (AS3): Scene 1 represents the scenario with signals incident at [10°, 13°, 16°], while Scene 2 represents the scenario with signals incident at [10°, 20°, 30°].
-In Scene 1, the angular interval between incident signals is smaller, which imposes higher demands on the model’s angular super-resolution capability.
-In general, since smaller angular intervals result in more similar steer vectors, algorithms based on linear space and compress sensing will perform better in Scene 2 than in Scene 1.
-On the other hand, deep learning algorithms are fundamentally designed to minimize the loss function.
-By minimizing the loss function across the entire training set, deep learning algorithms may achieve improved performance at smaller angular intervals.
-[…]
-RMSE $\boldsymbol{vs.}$ snapshots: To further demonstrate the model’s performance under varying snapshots (AS1) and angular ranges (AS3) for a ULA, we conducted simulation experiments in different snapshot scenarios.
-Similar to the previous experiment, we empirically evaluate the RMSE by averaging over 5,000 Monte Carlo trials.
-Two distinct scenarios are considered to investigate the model’s performance under different angular ranges.
-Scene 1 represents the scenario with signals incident at [5.1°, 30.3°, 54.5°], while Scene 2 represents the scenario with signals incident at [-9.9°, 15.3°, 39.5°].
-The demands on the model’s angular super-resolution are alleviated and the angular intervals are kept the same in two scenes.
-Thus, the larger angular range in Scene 2 leads to a lower angular resolution, which has a negative impact on the performance of the algorithms.
+7.
+within the high-precision regime ($0^{\circ}-3^{\circ}$), the cumulative probability of the proposed model is slightly lower than that of Learning-SPICE, SPE-CNN and MUSIC, while exhibiting advantage over the SubspaceNet, Unity-ESPRIT, and other algorithms.
+However, the probability of absolute errors exceeding $10^{\circ}$ is substantially reduced.
+This lower frequency of high-error occurrences highlights the model’s superior reliability and its ability to suppress outliers in challenging environments.
+Figure 8: RMSE $vs.$ SNR results under Scenarios 2 with Uniform configuration and Params [1].
+Figure 9: RMSE $vs.$ SNR results under Scenarios 3 with Uniform configuration and Params [1].
+To rigorously evaluate the model’s performance, we move beyond traditional sorting-based pairing schemes, which fail to accommodate cardinality mismatches and introduce systematic bias by discarding samples with missed detections.
+Instead, we utilize the Hungarian algorithm for optimal source association, with estimation error capped at $30^{\circ}$ to prevent extreme outliers from distorting the statistical results.
+Evaluation is conducted across both ’Raw’ and ’Matched’ configurations using a comprehensive suite of metrics: OSPA [46] (to jointly assess cardinality and localization accuracy), MAE, RMSE, Accuracy (Acc., success ratio within a $10^{\circ}$ tolerance), and the 10th/90th percentiles of the ECDF.
+These quantiles respectively characterize the high-precision capability and the robustness boundary of the model.
+The performance results for Scenario 1 (Uniform configuration, Params [1]) are detailed in Table II.
+In the table, algorithms with non-zero miss probability are shaded in gray, signifying a potential risk of target loss.
+Directional arrows ($\downarrow$ / $\uparrow$) are employed to indicate whether a higher or lower value represents superior performance.
+For clarity, the optimal values for each metric are bolded, while the second-best results are underlined.
+It is observed that the proposed algorithm consistently achieves the best or second-best performance across the majority of metrics, demonstrating its competitive edge in both estimation accuracy and reliability.
+V-B2 Simulation results in Scenario 2 and 3
+(a) TransDOA under Scene 1
+(b) MUSIC under Scene 1
+(c) TransDOA under Scene 2
+(d) MUSIC under Scene 2
+Figure 10: Scatter plots of our proposed model and MUSIC algorithm under Scenarios 4 with Deterministic configuration and Params [1].
+To evaluate the model’s performance in handling increased source cardinality and larger array apertures, we conduct further experiments under Scenarios 2 and 3.
+The RMSE $vs.$ SNR results under Scenario 2 is illustrated in Fig.
+9.
+It can be observed that the proposed approach achieves superior statistical performance in low SNR and limited snapshots scenarios compared with all the subspace-based methods, compressive sensing techniques, and state of the art deep learning algorithms.
+This indicates that the proposed model maintains robustness even in high-density source environment.
+Furthermore, the algorithm is evaluated under Scenario 3.
 As shown in Fig.
-10, a clear trend can be observed: for any given algorithm, the performance in Scene 1 is inferior to that in Scene 2, which can be attributed to the more challenging conditions in Scene 1.
-Furthermore, the proposed DOA-ViT model generally outperforms the other algorithms in this low SNR scenario.
-In Scene 1, our model achieves the best performance, surpassing the results achieved by MUSIC, $\ell_{1}$-SVD, Learning-SPICE, and Unity-ESPRIT.
-In Scene 2, the proposed model also attains the best performance, slightly outperforming the Learning-SPICE algorithm.
-These results highlight the superior effectiveness of our approach.
-V-B2 Simulation results in Scenario 2 and 3
-(a) ViT under Scene 1
-(b) MUSIC under Scene 1
-(c) ViT under Scene 2
-(d) MUSIC under Scene 2
-Figure 11: Scatter plots of our proposed model and MUSIC algorithm under Scenarios 4 with $snap=50$ and $SNR=-5db$.
-Scene 1 is configured with $\boldsymbol{\theta}=[-120,-60,0,60,120]$ and the corresponding azimuth angles $\boldsymbol{\phi}=[30,10,30,10,35]$.
-Scene 2 is configured with $\boldsymbol{\theta}=[-135,-120,-60,0,60]$, $\boldsymbol{\phi}=[35,30,10,30,10]$.
+9, the proposed model outperforms other algorithms in low SNR scenarios, except for cases where SNR=5dbSNR=-5db or $SNR=5db$.
+Due to the limitations of data-driven algorithm fitting accuracy, the proposed model exhibits a performance plateau in the range of 0 dB to 5 dB.
+Consequently, its performance is inferior to Unity-ESPRIT and $\ell_{1}$-SVD algorithms under the $SNR=5db$ scenario, as these analytical methods benefit more effectively from the increased number of array elements.
+The performance of the proposed model in DOA estimation is difficult to improve in high-precision scenarios, which is an area for future improvement.
+V-B3 Simulation results in Scenario 4
+To evaluate the model’s ability in two-dimensional scenarios, the proposed architecture is extended to jointly regress the values of elevation $\boldsymbol{\theta}$ and azimuth $\boldsymbol{\phi}$ under Scenario 4.
+Similar to one-dimensional DOA estimation, two-dimensional DOA estimation is inherently dependent on multiple factors.
+Therefore, we consider two DOA configurations with different angular separations.
+To facilitate a controlled comparison, we only modify one of the angular separation parameters between the two setups.
+Configuration 1 consists of incident signals at $\boldsymbol{\theta}=[-120,-60,0,60,120]$ with corresponding azimuth angles given by $\boldsymbol{\phi}=[30,10,30,10,35]$.
+In Configuration 2, the fifth target is relocated to test the resolution limits, resulting in $\boldsymbol{\theta}=[-135,-120,-60,0,60]$ and $\boldsymbol{\phi}=[35,30,10,30,10]$.
+A representative visualization of the results is presented in Fig.
+10, the symbol ”+” represents the true location of the target, while the dot in the figure represents the predicted value of the algorithm.
+Each predicted value shares the same color as its corresponding ground truth for clarity.
+Based on 100 Monte Carlo trials per configuration, it can be observed in Configuration 1 that the predicted results of the proposed model are predominantly clustered around the true DOAs.
+In contrast, while the MUSIC algorithm performs adequately for widely separated targets like $\theta=[-120,120]$ and $\phi=[30,35]$, but it tends to confuse the other three points, leading to degraded precision.
+In Configuration 2, the fifth target is moved adjacent to the first.
+It can be observed that the MUSIC algorithm fails to distinguish between the two targets, producing a single merged spectral peak.
+In contrast, the proposed model maintains high resolution, with predicted values still well clustered around their respective ground truth positions.
+These results highlight the superior capability of the model in achieving fine-grained localization in complex 2D DOA estimation tasks.
 (a) CDF of $\boldsymbol{\theta}$
 (b) CDF of $\boldsymbol{\phi}$
-Figure 12: The CDF of the absolute error under Scenarios 4 with $snap=50$ and $SNR=-5db$ is shown for: (a) $\boldsymbol{\theta}$, and (b) $\boldsymbol{\phi}$.
-The angles of the incident signals are randomly generated.
-To evaluate the model’s performance in handling a larger number of incident angles or with more array elements, we extend the testing from Scenario 1 to Scenario 2 and Scenario 3.
-We plotted the RMSE $vs.$ SNR results under Scenario 2 with $snap=10$ in Fig.
-10.
-In this case, the incident signals are randomly generated within the range of $[-60^{\circ},60^{\circ}]$, and the minimum angular interval is set to 3 degrees.
-If the generated angular interval is less than 3 degrees, the incident angle will be regenerated.
-Notably, it can be observed that, compared with all the subspace-based methods, compressive sensing methods, and the state of the art deep learning algorithms, the proposed approach achieves superior statistical performance in low SNR and small snapshots scenarios.
-This indicates that the proposed model can still deliver relatively robust performance even in scenarios involving a greater number of incident angles.
-Furthermore, the algorithm is tested under Scenario 3 with $\text{snap}=10$.
-In Fig.
-10, it can be observed that the proposed model outperforms other algorithms in low SNR and small snapshot scenarios, except for cases where $SNR=-5db$ or $SNR=5db$.
-$\ell_{1}$-SVD and Unity-ESPRIT algorithms demonstrates high accuracy in scenarios with a larger number of array elements.
-Due to the limitations of data-driven algorithm fitting accuracy, the proposed model could not show significant improvement in algorithm performance in the range of 0 dB to 5 dB.
-As a result, its performance is inferior to Unity-ESPRIT and $\ell_{1}$-SVD algorithms under the $SNR=5db$ scenario.
-The performance of the proposed model in DOA estimation is difficult to improve in high-precision and higher SNR scenarios, which is an area for future improvement.
-V-B3 Simulation results in Scenario 4
-To evaluate the model’s ability to extend from one-dimensional to two-dimensional scenarios, we expand the model to two dimensions by simultaneously regressing the values of $\theta$ and $\phi$, and model’s performance is tested in Scenario 4.
-Similar to one-dimensional DOA estimation, two-dimensional DOA estimation is simultaneously influenced by multiple parameters.
-Therefore, we consider two Scenes with different angular separations (AS3).
-Since it is uncommon to simultaneously vary the angular separations of all signals in practice scenarios, we only modify one of the angular separation parameters across the two scenes.
-In Scene 1, the incident signals are generated with $\boldsymbol{\theta}=[-120,-60,0,60,120]$, with the corresponding azimuth angles given by $\boldsymbol{\phi}=[30,10,30,10,35]$.
-Scene 2 is configured with $\boldsymbol{\theta}=[-135,-120,-60,0,60]$ and $\boldsymbol{\phi}=[35,30,10,30,10]$.
-A representative visualization of the results is presented in Fig.
-11, which facilitates an intuitive comparison of the algorithm’s performance.
-In Fig.
-11, the symbol ”+” represents the true location of the target, while the dot in the figure represents the predicted value of the algorithm.
-Additionally, the model’s predicted values and their corresponding true values are marked with the same color.
-100 Monte Carlo simulations were conducted for each scene.
-According to the experimental results in Scene 1, it can be observed that the predicted results of the proposed model are predominantly clustered around the true values.
-In contrast, the MUSIC algorithm achieves better DOA estimation performance at the more widely separated targets, like $\theta=[-120,120]$ and $\phi=[30,35]$, but it tends to confuse the other three points, resulting in poorer prediction performance.
-In Scene 2, we moved the fifth target from $\theta=[120]$ $\phi=[35]$ to $\theta=[-135]$ $\phi=[35]$, which is adjacent to the first target.
-It can be observed that the MUSIC algorithm fails to distinguish between the two targets, forming a single spectral peak between them, whereas the predicted results of the proposed model remain scattered around the true values.
+Figure 11: The CDF of the absolute error under Scenarios 4 with $snap=50$ and SNR=5dbSNR=-5db is shown for: (a) $\boldsymbol{\theta}$, and (b) $\boldsymbol{\phi}$.
 (a) $\rho=0$
@@
 (c) $\rho=1$
-Figure 13: Three different levels of array imperfections are proposed to evaluate the performance of our proposed model and traditional algorithms: (a) $\rho=0$, no array imperfections, (b) $\rho=0.5$, moderate array imperfections, and (c) $\rho=1$, severe array imperfections.
-Figure 14: RMSE $vs.$ SNR results under Scenarios 1 with $snap=10$ and array imperfections of $\rho=1$.
-The transfer learning approach is employed to enhance the model’s performance under array imperfection scenarios.
+Figure 12: RMSE $vs.$ SNR results under Scenarios 1 with Uniform configuration and Params [2].
+Three different levels of array imperfections are proposed: (a) $\rho=0$, no array imperfections, (b) $\rho=0.5$, moderate array imperfections, and (c) $\rho=1$, severe array imperfections.
+Figure 13: RMSE $vs.$ SNR results under Scenarios 1 with Uniform configuration, Params [2] and $\rho=1$.
 The dashed line represents the case where only 50 samples are used for transferring learning, while the solid line indicates the use of 500 samples.
-Furthermore, 1000 simulations are performed under Scenario 4 with $snap=50$ and $SNR=-5\ \mathrm{dB}$ to evaluate the statistical performance of the proposed model, where the angles of the incident signals are randomly generated.
-The CDF of the proposed model is presented in Fig.
-12.
-It can be observed that the MUSIC algorithm performs better on CDF(10).
-However, its performance deteriorates on CDF(30).
-This suggests that while the MUSIC algorithm is likely to achieve more accurate DOA estimations for certain targets, it also shows a higher probability of producing significant DOA estimation errors.
-In contrast, the proposed model demonstrates a more consistent overall stability in its predictions.
-This experiment also demonstrates that the CDF of Absolute Error serves as an effective metric in DOA estimation tasks.
-Compared with RMSE, it provides a more comprehensive evaluation of the model’s performance.
-Supported by the Cramér-Rao bound as a theoretical benchmark, RMSE offers a significant reference value, whereas the CDF of Absolute Error achieves a more thorough assessment of the model’s capabilities.
-Through the simulation results, it is evident that the proposed model achieves a greater advantage in terms of RMSE.
-However, through the analysis of the CDF of Absolute Error, it has certain limitations results in the precision DOA estimation.
+Furthermore, 1000 simulations are performed under Scenario 4 with Uniform configuration and Params [1].
+The ECDF of the absolute DOA estimation errors is presented in Fig.
+11.
+It can be observed that the MUSIC algorithm performs better on $\text{CDF}(5^{\circ})$.
+However, its performance deteriorates on $\text{CDF}(15^{\circ})$.
+This suggests that although MUSIC can achieve high-precision estimations for certain samples, it is more prone to large estimation outliers.
+In contrast, the proposed model exhibits superior consistency and robustness.
+This experiment further validates the effectiveness of ECDF as a comprehensive metric for DOA estimation.
+While RMSE remains a critical benchmark, especially when referenced against the CRLB, it only captures the average performance.
+In comparison, the ECDF provides a more detailed perspective on the error distribution.
+The results demonstrate that while the proposed model holds a clear advantage in terms of RMSE, the ECDF analysis reveals its subtle limitations in achieving high precision compared to MUSIC’s best-case scenarios.
 V-C Transfer Learning Approach for Array imperfection
 The presence of array imperfections introduces inconsistencies between the received signal model and the ideal model, leading to a degradation in the DOA estimation performance of the model.
-Scenarios involving multiple array imperfections, particularly under low SNR and small snapshot conditions, are inherently complex and challenging to handle.
+Scenarios involving multiple array imperfections, particularly under low SNR and limited snapshot conditions, are inherently complex and challenging to handle.
 In this section, we demonstrate the effectiveness of the proposed transfer learning approach in these scenarios through simulation experiments.
 V-C1 Transfer Learning in Scenario 1
-Figure 15: DOA estimation under Scenarios 1 with $SNR=5db$ and $snap=10$, $\rho=1$.
-Three incident signals are generated and the interval is set to $\Delta\hat{\theta}={10}^{\circ}$.
-The line represents the true location of the target, while the dot in the figure represents the predicted value of the algorithm.
-Additionally, the model’s predicted values and their corresponding true values are marked with the same color.
-TABLE II: Transfer Learning Results with Varying Sample Number
+Figure 14: DOA estimation under Scenarios 1 with $SNR=5db$, $snap=10$ and $\rho=1$.
+The interval is set to $\Delta\hat{\theta}={10}^{\circ}$.
+The line represents the true location of the target, while the dot represents the predicted value of the algorithm.
+TABLE III: Transfer Learning Results with Varying Sample Number
 Sample Number 20 50 100 200 300 500 800 1000
@@
 Trnasfer Learning 13.66 13.27 13.06 12.72 12.60 12.54 12.21 11.48
-(a) ViT without transfer learning
-(b) ViT with transfer learning
-Figure 16: DOA estimation under Scenarios 1 with $SNR=5db$, $snap=10$ and array imperfections of $\rho=1$.
+(a) TransDOA without transfer learning
+(b) TransDOA with transfer learning
+Figure 15: DOA estimation under Scenarios 1 with Equidistant configuration, $SNR=5db$, $snap=10$ and $\rho=1$.
 100 Monte-Carlo simulations are conducted and the incident signal is configured with $\boldsymbol{\theta}=[-135,-120,-60,0,60]$, $\boldsymbol{\phi}=[35,30,10,30,10]$.
-To demonstrate the performance degradation of the deep learning model under array imperfections, we illustrate the model’s performance under scenarios with multiple array imperfections, as depicted in Fig.
-13.
-The position bias, amplitude and phase inconsistency, and mutual coupling effects of the array systems are simultaneously considered, and the intensity of these array imperfections is controlled by a hyperparameter $\rho$.
+To demonstrate the performance degradation of the deep learning model under array imperfections, we illustrate the model’s performance under scenarios with multiple array imperfections in Fig.
+12.
+These imperfections include position bias, amplitude and phase inconsistency, and mutual coupling effects, with their overall severity controlled by a hyperparameter $\rho$.
 As shown in Fig.
-13, It can be observed that under the influence of multiple array imperfections, the statistical performance of deep learning-based methods (the proposed model, Learning-SPICE and SPE-CNN), compressed sensing-based methods ($\ell_{1}$-SVD), and subspace-based methods (Unity-ESPRIT and MUSIC) all deteriorates noticeably.
-Although the proposed model achieves optimal performance under the ideal signal model, particularly in low SNR and small snapshots scenarios, its performance significantly degrades when subjected to array imperfections.
+12, It can be observed that under the influence of multiple array imperfections, the statistical performance of deep learning-based methods (the proposed model, Learning-SPICE and SPE-CNN), compressed sensing-based methods ($\ell_{1}$-SVD), and subspace-based methods (Unity-ESPRIT and MUSIC) all deteriorates noticeably.
+Although the proposed model achieves optimal performance under the ideal signal model, particularly in low SNR and limited snapshots scenarios, its performance significantly degrades when subjected to array imperfections.
 This degradation becomes particularly pronounced in scenarios with relatively higher SNR.
 When $\rho=1$, the performance of the proposed model declines level similar to the traditional algorithms at $\rho=0$.
-To validate the effectiveness of the transfer learning approach, we evaluate the performance of the model adapted through transfer learning under varying SNR through 5,000 Monte Carlo simulations.
+The efficacy of the transfer learning scheme is further validated through 5,000 Monte Carlo trials across varying SNR levels.
 As shown in Fig.
-14, It can be observed that the proposed model achieves significant performance even with only 50 samples, demonstrating accurate DOA estimation under low SNR and small snapshot scenarios.
+13, the proposed model achieves significant performance even with only 50 samples, demonstrating accurate DOA estimation under low SNR and limited snapshot scenarios.
 Moreover, when more samples (500 samples) are used, the model’s performance is further enhanced.
-To demonstrate the superiority of the proposed transfer learning algorithm, a comparison experiment is carried out under Scenario 1 with $snap=10$, $SNR=-5$, and array imperfections $\rho=1$.
-Our method is compared with two other approaches: (1) directly training the model through the samples with array imperfections (Directly Train), and (2) fine-tuning the model trained on simulated data using samples with array imperfections (Fine-tune).
-The performance of these three methods are evaluated across varying numbers of array imperfection samples, as shown in Table II.
+To demonstrate the superiority of the proposed transfer learning algorithm, a comparison experiment is carried out under Scenario 1 with $snap=10$, SNR=5SNR=-5 (Params [1]), and array imperfections $\rho=1$.
+Our method is compared with two other approaches: (1) Direct Train, which involves training from scratch on imperfect data, and (2) Fine-tune, where a pre-trained model on simulated data is adapted samples with array imperfections.
+The performance of these three methods are evaluated across a varying number of training samples, as summarized in Table III.
 The results indicate that the proposed transfer learning algorithm consistently achieves superior performance in most cases.
 This demonstrates that a well-designed transfer learning strategy can utilize the information of limited samples more effectively, outperforming both the Fine-tune and Directly Train methods.
-To illustrate the performance of the proposed model after transfer learning, an experiment is conducted to visually demonstrate our model’s performance, and the results are presented in Fig.
-15.
-In this experiment, three incident angles are generated with a constant interval of $\Delta\hat{\theta}={10}^{\circ}$, covering the range from -60 to 60 degrees, with the incident angles linearly scaled with a step of 0.5 degrees.
-From the experimental results, it can be observed that the proposed model exhibits significant bias before applying transfer learning, particularly when the incident angles are close to -60 degrees.
-However, after applying transfer learning, the predict results of the proposed model become concentrated around the true values, demonstrating high DOA estimation accuracy.
-In contrast, due to the combined effects of multiple array imperfections, the performance of the MUSIC and $\ell_{1}$-SVD algorithm deteriorates, exhibiting substantial deviations.
-Meanwhile, Unity-ESPRIT and SPE-CNN show significant deviations when the incident angles span a wider range.
-SubspaceNet exhibits noticeable distortions, failing to achieve accurate DOA estimation.
-Learning-SPICE, SPE-CNN and ASL-2 demonstrate enhanced resilience to array imperfections; however, systematic distortion remains observable in the predict results.
+The impact of transfer learning is further highlighted in Fig.
+14, considering the Equidistant configuration of Scenario 1 with $SNR=5db$, $snap=10$.
+In this experiment, three DOAs are generated with a constant interval of $\Delta\hat{\theta}={10}^{\circ}$, covering the range from -60 to 60 degrees.
+Evidently, the proposed model exhibits significant bias before applying transfer learning, particularly when the DOAs are close to -60 degrees.
+Transfer learning effectively rectifies these estimation errors, forcing the predicted estimates to converge tightly around the ground truth.
+In contrast, the performance of the MUSIC, $\ell_{1}$-SVD and SubspaceNet suffer from substantial deviations due to array imperfections.
+Meanwhile, Unity-ESPRIT and ASL-2 fail to maintain reliable estimation as the angular range expands.
+Learning-SPICE and SPE-CNNASL-2 demonstrate enhanced resilience to array imperfections; however, systematic distortion remains in the predict results.
 V-C2 Transfer Learning in Scenario 4
-To further validate the effectiveness of the proposed transfer learning algorithm, we extend the algorithm from one-dimensional to two-dimensional.
-The 2D DOA estimation results of the ViT before and after applying transfer learning are shown in Fig.
-16.
-It can be clearly observed that prior to the application of the transfer learning algorithm, the model fails to effectively distinguish between the first target ($\theta=[-135]$, $\phi=[35]$) and the second target ($\theta=[-120]$, $\phi=[30]$) and exhibits significant bias for orther targets due to array imperfections.
+To further validate the scalability of the transfer learning approach, we extend its application from one-dimensional to two-dimensional.
+Fig.
+15 presents the DOA estimation results of the TransDOA both before and after applying transfer learning.
+It can be clearly observed that prior to the application of the transfer learning algorithm, the model fails to effectively distinguish between the first two targets ($\theta=[-135]$, $\phi=[35]$ and $\theta=[-120]$, $\phi=[30]$) and suffers from significant bias for other targets due to array imperfections.
 However, after the transfer learning algorithm is applied, the prediction bias was substantially corrected, and the predicted values were able to converge near the true target.
 VI CONCLUSION
-In this work, we develop a DOA-ViT model for DOA estimation, which leverages the attention mechanism to capture global dependencies.
+In this work, we develop a TransDOA model for DOA estimation, which leverages the attention mechanism to capture global dependencies.
 By directly learning the mapping from the sampled SCM to DOA values, the proposed method can be easily extended to arbitrary array geometries or 2D DOA estimation scenarios, demonstrating strong adaptability.
-Extensive simulations are conducted to evaluate the performance and robustness of the proposed model under various conditions: i) different SNRs and numbers of snapshots, ii) varying numbers of array elements $M$ and incident signals $K$, and iii) different angular separations and signal ranges.
+Extensive simulations are conducted to evaluate the performance and robustness of the proposed model.
 Furthermore, deep learning based DOA estimators often suffer from severe performance degradation in practical scenarios with array imperfections.
 To overcome this limitation, we propose a supervised transfer learning strategy to mitigate the performance degradation in practical scenarios.
-By aligning the features of the source and target domains in a supervised manner, the proposed method demonstrates improved results in scenarios with array imperfections.
-It is worth noticing that this work is the first to introduce an advanced transfer learning strategy to address array imperfections, rather than simply employing fine-tune.
+By aligning the features of the source and target domains in a supervised manner, the proposed method demonstrates improved results in practical scenarios with array imperfections.
+It is worth noticing that this work is the first to introduce an advanced transfer learning strategy to address array imperfections.
 Investigating how to effectively leverage transfer learning methods to improve the performance of deep learning models in scenarios with array imperfections is a significant and worthwhile research direction, and has the potential to enhance the performance of deep learning algorithms in practical applications.